### v2

# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

Abstract
In this paper, we investigate how structural properties of the constraint system impact the oracle complexity of smooth non-convex optimization problems with convex inequality constraints over a simple polytope. In particular, we show that, under a local error bound condition with exponent $d\in[1,2]$ on constraint functions, an inexact Moreau envelope Lagrangian method can attain an $\epsilon$-Karush–Kuhn–Tucker point with $\tilde{O}(\epsilon^{-2d})$ gradient oracle complexity. When $d=1$, this result matches the best-known complexity in literature up to logarithmic factors. Importantly, the assumed error bound condition with any $d\in[1,2]$ is strictly weaker than the local linear independence constraint qualification that is required to achieve the best-known complexity. Our results clarify the interplay between error bound conditions of constraints and algorithmic complexity, and extend complexity guarantees to a broader class of constrained non-convex problems.
1 Introduction
In this paper, we study first-order methods (FOMs) for nonlinear constrained optimization problems formulated as
$\min_{{\mathbf{x}}\in\mathcal{X}}f({\mathbf{x}})\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ {\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},$ (1)
where $\mathcal{X}\subset\mathbb{R}^{n}$ is a bounded polytope that allows a computationally easy projection operator, $f$ is smooth but not necessarily convex, and $g_{i}$ for $i=1,\dots,m$ are smooth and convex. Since (1) is non-convex, computing its optimal or $\epsilon$-optimal solution is in general intractable. Therefore, the objective of this study is to compute an $\epsilon$-Karush–Kuhn–Tucker ($\epsilon$-KKT) point of (1) (see Definition 3.2) using an FOM.
The FOM studied in this paper is the inexact Moreau envelope Lagrangian (iMELa) method (see Algorithm 1). This method is related to the proximal augmented Lagrangian method (ALM), a technique with a long history [12, 39, 43, 42, 44, 45] that remains an active area of research [28, 19, 20, 34, 22, 21, 53, 54, 55, 40, 52]. Following the literature, we measure the computational efficiency of the iMELa method by its first-order oracle complexity, which is the number of gradient evaluations of $f$ or $g_{i}$ for $i=1,\dots,m$ required to reach an $\epsilon$-KKT point.
[…]
Beyond the complexity results, this work contributes to a deeper understanding of the interaction between the convergence behavior of an FOM and EBCs of constraint functions. Error bound conditions have a long history in optimization, due to their connections to constraint qualifications and their impact on the convergence of optimization algorithms (see, e.g., Lewis and Pang [24], Luo and Pang [32], Pang [38], Li [25]). While EBCs on objective functions have been extensively studied and exploited to accelerate the convergence of FOMs, the role of EBCs on constraint functions has received relatively limited attention. This work helps bridge this gap by showing how such conditions are utilized in the convergence analysis of an FOM for nonconvex constrained optimization.
2 Related Works
In the recent literature [3, 33, 17, 16, 30], subgradient-based methods have been developed for non-smooth continuous optimization problems with convex or weakly convex inequality constraints and achieved the oracle complexity of $O(\epsilon^{-4})$ for finding a nearly $\epsilon$-KKT point. Although those methods can be applied to (1), their analysis does not utilize the smoothness of the functions and thus yields a complexity higher than the one in this work when the constraints are convex.
The iMELa method we study is closely related to the classical augmented Lagrangian method (ALM), which has a long history of study [12, 39, 43, 42, 44, 45] and is still one of the most effective approaches for constrained optimization. Below, we review the recent studies on ALM and its variants, including the penalty methods, for the following general problem
$\min_{{\mathbf{x}}}$ $\penalty 10000\ f({\mathbf{x}})+r({\mathbf{x}})$ (2)
s.t. $\penalty 10000\ {\mathbf{g}}({\mathbf{x}})=(g_{1}({\mathbf{x}}),\dots,g_{m}({\mathbf{x}}))^{\top}\leq{\mathbf{0}},$
$\penalty 10000\ {\mathbf{h}}({\mathbf{x}})=(h_{1}({\mathbf{x}}),\dots,h_{l}({\mathbf{x}}))^{\top}={\mathbf{0}},$
where $f$, $g_{i}$, $i=1,\dots,m$, and $h_{j}$, $j=1,\dots,l$, are smooth and have Lipschitz continuous gradients, and $r$ is a proper convex lower semi-continuous function that allows a computationally easy proximal mapping. When (2) is convex, the oracle complexity of the ALM for finding an $\epsilon$-optimal solution is well studied. See, e.g. Lan and Monteiro [23], Xu [50, 49], He and Yuan [11], Liu et al. [31] for the results under different settings. We next focus on the works that are applicable to (2) or its special cases when the problem is non-convex.
In several studies on ALM (e.g. Hong [14], Hajinezhad and Hong [10], Zeng et al. [52]), the efficiency of an algorithm is characterized by its iteration complexity, defined as the number of main iterations needed for computing an $\epsilon$-KKT point. Since a (proximal) augmented Lagrangian subproblem must be solved in each main iteration of ALM—typically using a separate FOM—the oracle complexity is generally higher than the iteration complexity.
Methods for linear constraints. Suppose $g_{i}$’s are not present and $h_{j}$’s are linear in (2). When $r\equiv 0$, Hong [14] introduced a proximal primal-dual algorithm (prox-PDA) that finds an $\epsilon$-KKT point of (2) with an iteration complexity of $O(\epsilon^{-2})$. When $r$ in (2) is the characteristic function of a box or a bounded polytope, it has been shown in Zhang and Luo [53, 54] that a smoothed proximal ALM (SP-ALM), which generalizes the classical proximal ALM [45], can achieve an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. When $r$ is the characteristic function of a compact set defined by convex inequalities, it has been shown that SP-ALM also achieves an $O(\epsilon^{-2})$ oracle complexity under CRCQ [55].
Suppose no additional structural assumption is made on $r$. Hajinezhad and Hong [10] developed a variant of prox-PDA with an iteration complexity of $O(\epsilon^{-4})$. Zeng et al. [52] proposed the Moreau envelope ALM (MEAL), a type of proximal ALM, which achieves an iteration complexity of $o(\epsilon^{-2})$ when $f$ satisfies an implicit Lipschitz subgradient property and $O(\epsilon^{-2})$ when $f$ satisfies an implicit bounded subgradient property. The convergence property of MEAL is also established when the augmented Lagrangian function satisfies the Kurdyka–Łojasiewicz property [52]. Kong et al. [19, 20] proposed a quadratic penalty accelerated inexact proximal point (QP-AIPP) method, which finds an $\epsilon$-KKT point with an oracle complexity of $\tilde{O}(\epsilon^{-3})$. Under additional mildly strong assumptions, a reduced oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ is achieved by an inexact ALM (iALM) [27], an inexact proximal accelerated augmented Lagrangian (IPAAL) method [34], and an inner accelerated inexact proximal augmented Lagrangian (IAIPAL) method [22].
Methods for nonlinear convex constraints. Suppose $g_{i}$’s are convex and $h_{j}$’s are linear in (2). Lin et al. [28] proposed an inexact proximal point penalty (iPPP) method that achieves an oracle complexity of $\tilde{O}(\epsilon^{-2.5})$ under a Slater’s condition. Under the similar assumptions, Li and Xu [26] extended this approach to a hybrid method combining ALM and the penalty method with the same complexity. Dahal et al. [8] provided a damped proximal ALM that achieves the same order of oracle complexity. The iALM by Li et al. [27] can also achieve the same complexity but requires an additional regularity assumption (see (3) below). The IAIPAL introduced in Kong et al. [22] was also extended by Kong et al. [21] to handle nonlinear convex constraints and obtains an $\tilde{O}(\epsilon^{-3})$ oracle complexity.
When $r$ is the characteristic function of a compact set defined by convex inequalities, Pu et al. [40] developed a smoothed proximal Lagrangian method (SP-LM), which finds an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-2})$. In contrast, our method assumes $r$ is the characteristic function of a compact polytope and requires a complexity of $\tilde{O}(\epsilon^{-2})$. Besides the assumption on $r$, there are two key differences between their method and ours in the updating schemes. First, their method projects the dual variables onto a compact, artificially constructed set, whereas our method only needs to project the dual variables onto $\mathbb{R}^{m}_{+}$. Second, their method only uses a single loop but ours employs a double-loop structure. More importantly, their method relies on a Slater-like condition and a local LICQ condition at all KKT points, whereas we assume a Slater’s condition and a local error bound condition on each active subset of the constraint set (see Assumption 3.3). We will show (see Proposition 3.4) that, if the local LICQ condition in Pu et al. [40] holds over all KKT points, our local error bound assumption also holds. However, the converse is not true. Thus, our work extends the complexity analysis of SP-LM to the case not covered in Pu et al. [40].
Methods for nonconvex constraints. Suppose $g_{i}$’s are non-convex or $h_{j}$’s are nonlinear in (2). In this case, even finding a feasible solution for (2) is generally intractable. When $r$ is the characteristic function of a box and ${\mathbf{g}}\equiv{\mathbf{0}}$, Curtis et al. [6, 7] proposed an augmented Lagrangian trust-region method with the global asymptotic convergence guarantee. Their methods may produce an infeasible stationary point due to the non-convex constraints. In general, additional assumptions are needed to obtain a feasible stationary point. For example, assuming a (nearly) feasible solution can be easily obtained, the iPPP method [28] and a scaled dual descent alternating direction method of multipliers (SDD-ADMM) proposed by Sun and Sun [48] can find an $\epsilon$-KKT point with an oracle complexity of $O(\epsilon^{-4})$. Another common assumption made in literature is the regularity condition, which assumes that there exists $\zeta>0$ such that
$\zeta\sqrt{\|[{\mathbf{g}}({\mathbf{x}})]_{+}\|^{2}+\|{\mathbf{h}}({\mathbf{x}})\|^{2}}\leq\text{dist}\left(-\nabla{\mathbf{g}}({\mathbf{x}})[{\mathbf{g}}({\mathbf{x}})]_{+}-\nabla{\mathbf{h}}({\mathbf{x}}){\mathbf{h}}({\mathbf{x}}),\partial r({\mathbf{x}})\right)$ (3)
for any ${\mathbf{x}}\in\text{dom}(r)$. When (3) holds on the constraints and $g_{i}$’s are not present, Sahin et al. [47] showed that iALM can find an $\epsilon$-KKT point with a complexity of $\tilde{O}(\epsilon^{-4})$. This complexity is reduced to $\tilde{O}(\epsilon^{-3})$ by the iALM [27] [^2] Although the problem considered in Li et al. [27] involves only equality constraints, their results can be extended to the general problem (2), as noted in Li et al. [27, Remark 6]. and the iPPP method [28] with the inequality constraints, and to $O(\epsilon^{-3})$ by the SDD-ADMM method [48] without the inequality constraints when there is only one block in $\mathbf{x}$}. Moreover, the SDD-ADMM method [48] can achieve a complexity of $O(\epsilon^{-2})$ under addtional stronger assumptions.
3 Preliminaries
For vectors, $\|\cdot\|$ denotes the Euclidean ($\ell_{2}$) norm and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. For matrices, $\|\cdot\|$ denotes the spectral norm. For a differentiable function $h$ on $\mathbb{R}^{n}$, we say $h$ is $L$-smooth ($L\geq 0$) on $\mathcal{X}$ if $\|\nabla h({\mathbf{x}})-\nabla h({\mathbf{x}}^{\prime})\|\leq L\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$, $\mu$-strongly convex ($\mu\geq 0$) on $\mathcal{X}$ if $h({\mathbf{x}})\geq h({\mathbf{x}}^{\prime})+\langle\nabla h({\mathbf{x}}^{\prime}),{\mathbf{x}}-{\mathbf{x}}^{\prime}\rangle+\frac{\mu}{2}\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|^{2}$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$, and $\rho$-weakly convex ($\rho\geq 0$) on $\mathcal{X}$ if $h({\mathbf{x}})\geq h({\mathbf{x}}^{\prime})+\langle\nabla h({\mathbf{x}}^{\prime}),{\mathbf{x}}-{\mathbf{x}}^{\prime}\rangle-\frac{\rho}{2}\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|^{2}$ for any $({\mathbf{x}},{\mathbf{x}}^{\prime})\in\mathcal{X}\times\mathcal{X}$. Note that the $L$-smoothness of $h$ implies that $h$ is $L$-weakly convex but $h$ can have a smaller weak convexity constant.
For any ${\mathbf{x}}\in\mathbb{R}^{n}$ and $\delta>0$, we denote by $\mathbb{B}({\mathbf{x}},\delta)$ the Euclidean ball centered at $\mathbf{x}$} with radius $\delta$. For a closed convex set $\mathcal{B}$, we denote its normal cone at $\mathbf{x}$} by $\mathcal{N}_{\mathcal{B}}({\mathbf{x}})$. Let $\delta_{\mathcal{B}}({\mathbf{x}})$ be the zero-infinity characteristic function of $\mathcal{B}$, i.e., $\delta_{\mathcal{B}}({\mathbf{x}})=0$ if $\mathbf{x}$$\in\mathcal{B}$ and $\infty$ otherwise. Let $\text{proj}_{\mathcal{B}}(\cdot)$ be the projection mapping onto $\mathcal{B}$ and $\text{dist}({\mathbf{x}},\mathcal{A}):=\min_{{\mathbf{y}}\in\mathcal{A}}\|{\mathbf{x}}-{\mathbf{y}}\|$ for a nonempty closed set $\mathcal{A}$. Define $\cdot]_{+}:=\max\{\cdot,0\}$. Denote $m]=\{1,\dots,m\}$ and $l]=\{1,\dots,l\}$.
The following assumption is made throughout the paper.
Assumption 3.1.
The following statements hold:
  • A.
    $\mathcal{X}=\big\{{\mathbf{x}}\in\mathbb{R}^{n}:{\mathbf{A}}{\mathbf{x}}\leq{\mathbf{b}}\big\}$ with ${\mathbf{A}}\in\mathbb{R}^{l\times n}$ and ${\mathbf{b}}\in\mathbb{R}^{l}$ is compact with diameter $D_{\mathcal{X}}:=\max_{{\mathbf{x}},{\mathbf{x}}^{\prime}\in\mathcal{X}}\|{\mathbf{x}}-{\mathbf{x}}^{\prime}\|<+\infty$.
  • B.
    $\min_{{\mathbf{x}}\in\mathcal{X}}f({\mathbf{x}})\geq\underline{f}$ for some $\underline{f}\in\mathbb{R}$.
  • C.
    $f$ and $g_{i}$, $i\in[m]$, are $L$-smooth, and each $g_{i}$ is convex.
  • D.
    $\max_{{\mathbf{x}}\in\mathcal{X}}\{\|\nabla f({\mathbf{x}})\|\}\leq B_{f}$ for some $B_{f}\geq 0$.
  • E.
    $\max_{{\mathbf{x}}\in\mathcal{X}}\max\{\|{\mathbf{g}}({\mathbf{x}})\|,\|\nabla{\mathbf{g}}({\mathbf{x}})\|\}\leq B_{g}$ for some $B_{g}\geq 0$.
  • F.
    (Slater’s condition) There exists ${\mathbf{x}}_{\textup{feas}}\in\mathcal{X}$ such that $g_{i}({\mathbf{x}}_{\textup{feas}})<0$, $i\in[m]$.
We denote the Lagrangian function of (1) by
$\mathcal{L}({\mathbf{x}},\boldsymbol{\lambda}):=f({\mathbf{x}})+\textstyle\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}}),$ (4)
where $\boldsymbol{\lambda}=(\lambda_{1},\dots,\lambda_{m})^{\top}$ is a vector of Lagrangian multipliers. We consider a numerical method for finding an $\epsilon$-KKT point of (1) defined as follows.
Definition 3.2.
A point ${\mathbf{x}}\in\mathcal{X}$ is an $\epsilon$-KKT point of problem (1) if there exists $\boldsymbol{\lambda}\geq{\mathbf{0}}$ such that
$\textup{dist}\big(-\nabla f({\mathbf{x}})-\textstyle\sum_{i=1}^{m}\lambda_{i}\cdot\nabla g_{i}({\mathbf{x}}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}})\big)\leq\epsilon,$ (5)
$\|[{\mathbf{g}}({\mathbf{x}})]_{+}\|\leq\epsilon,$ (6)
$\textstyle\sum_{i=1}^{m}|\lambda_{i}g_{i}({\mathbf{x}})|\leq\epsilon.$ (7)
For any $I_{g}\subseteq[m]$ and $I_{A}\subseteq[l]$, we define an active feasible subset as
[…]
When $\mathcal{S}(I_{g},I_{A})\neq\emptyset$, we call $\mathcal{S}(I_{g},I_{A})$ the active feasible subset induced by $I_{g}$ and $I_{A}$. For any ${\mathbf{x}}\in\mathcal{X}$, let
$J_{g}({\mathbf{x}}):=\{i\in[m]:g_{i}({\mathbf{x}})=0\},\quad J_{A}({\mathbf{x}}):=\{j\in[l]:[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}=0\}$
be the index sets of the constraints active at $\mathbf{x}$}. Let $\mathcal{X}^{*}$ denote the set of KKT points of (1). We assume the constraints of (1) satisfy a uniform local EBC near each KKT point.
Assumption 3.3 (Local EBC with exponent $d$).
There exist $\delta>0$, $\gamma>0$, and $d\in[1,2]$ such that, for any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ and any ${\mathbf{x}}\in\mathbb{B}({\mathbf{x}}^{*},\delta)$, it holds for every index pair $I=(I_{g},I_{A})$ with $I_{g}\subseteq J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subseteq J_{A}({\mathbf{x}}^{*})$ that
[…]
Inequality (14) states that each active feasible set $\mathcal{S}(I_{g},I_{A})$ defined in (10) satisfies a local EBC with exponent $d\in[1,2]$, where the constants $\delta$ and $\gamma$ are uniform over all $\mathbf{x}$} sufficiently close to $\mathcal{X}^{*}$ and all admissible index pairs $(I_{g},I_{A})$. Similar EBCs have been studied in constrained optimization; see, e.g., Yang and Lin [51], Lin et al. [29]. We further show that Assumption 3.3 holds whenever the constraints in (1) satisfy the local linear independence constraint qualification (LICQ) assumed in Pu et al. [40, Assumption 2.2]. Specifically, the local LICQ requires that there exists $\zeta>0$ such that, for any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$,
$\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{J_{g}({\mathbf{x}}^{*})}({\mathbf{x}}^{*}),{\mathbf{A}}_{J_{A}({\mathbf{x}}^{*})}^{\top}\big]\big)\geq\zeta,$ (15)
where $\sigma_{\min}(\cdot)$ denotes the smallest singular value. Under (15), each set $\mathcal{S}(I_{g},I_{A})$ satisfies a local EBC with $d=1$, which in turn implies Assumption 3.3 for any $d\in[1,2]$. On the contrary, Assumption 3.3 does not imply the local LICQ and thus is strictly weaker. See the following proposition whose formal proof is provided in Appendix A.
Proposition 3.4.
If the constraints of (1) satisfy the local LICQ in (15), then Assumption 3.3 holds, but not vice versa.
Throughout the paper, we assume that $p$ is a constant satisfying $p>L$. Following Zhang and Luo [53, 54], Zhang et al. [55], given any ${\mathbf{z}}\in\mathcal{X}$, we define
$v({\mathbf{z}}):=$ $\penalty 10000\ \min_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},$ (16)
${\mathbf{x}}({\mathbf{z}}):=$ $\penalty 10000\ \mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X},\,g_{i}({\mathbf{x}})\leq 0,\,i\in[m]}\left\{f({\mathbf{x}})+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}\right\},$ (17)
and $\boldsymbol{\lambda}({\mathbf{z}})$ is the optimal Lagrangian multipliers corresponding to ${\mathbf{x}}({\mathbf{z}})$. Note that (16) is a convex optimization problem as $p>L$ and $v({\mathbf{z}})$ is differentiable by Danskin’s Theorem (see, e.g. Bertsekas [2, Proposition B.25]) with
$\nabla v({\mathbf{z}})=p({\mathbf{z}}-{\mathbf{x}}({\mathbf{z}})).$ (18)
Like (4), the Lagrangian function of (16) is
$\hskip-3.61371pt\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}):=f({\mathbf{x}})+{\textstyle\sum_{i=1}^{m}\lambda_{i}g_{i}({\mathbf{x}})}+\frac{p}{2}\|{\mathbf{x}}-{\mathbf{z}}\|^{2}.$ (19)
Similar to the notations in Zhang and Luo [53, 54], Zhang et al. [55], we define
$d(\boldsymbol{\lambda},{\mathbf{z}}):=$ $\penalty 10000\ \min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}),$ (20)
${\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}):=$ $\penalty 10000\ \mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}},\boldsymbol{\lambda}).$ (21)
By the strong duality (see, e.g. Rockafellar [41, Sec. 28–30]), it holds that
$v({\mathbf{z}})=\max_{\boldsymbol{\lambda}\geq{\mathbf{0}}}d(\boldsymbol{\lambda},{\mathbf{z}}).$ (22)
4 Inexact Moreau Envelope Lagrangian Method
We present the inexact Moreau envelope Lagrangian (iMELa) method, shown in Algorithm 1, for finding an $\epsilon$-KKT point of (1). At iteration $t$, the method first updates the dual variable $\boldsymbol{\lambda}^{(t+1)}$ via a projected gradient ascent step on function $\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda})$ at $\boldsymbol{\lambda}=\boldsymbol{\lambda}^{(t)}$ with step-size $\tau_{t}$. Then, given the strongly convex subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, we denote its optimal solution by
$\tilde{{\mathbf{x}}}^{(t+1)}:={\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$
in the rest of the paper for simplicity of notation. We approximately solve this subproblem by computing an $\epsilon_{t}$-stationary point ${\mathbf{x}}^{(t+1)}\in\mathcal{X}$ with $\epsilon_{t}\in(0,1)$, satisfying
$\text{dist}\big(-\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)\leq\epsilon_{t}.$ (23)
Such a point can be computed using the accelerated projected gradient (APG) method by Nesterov [37, Eq. (2.2.63)] with complexity provided shortly. Moreover, condition (23) can be numerically verified and thus used as a stopping criterion for the APG method, provided that $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})$ has a simple structure. Finally, the proximal center $\mathbf{z}$} is updated via a convex combination of ${\mathbf{z}}^{(t)}$ and ${\mathbf{x}}^{(t+1)}$ with a weight $\theta_{t}\in[0,1]$, ensuring ${\mathbf{z}}^{(t)}\in\mathcal{X}$ for all $t\geq 0$.
Algorithm 1 Inexact Moreau Envelope Lagrangian (iMELa) Method
1:Input: total number of iterations $T$, proximal parameter $p>0$, step-sizes $\{\tau_{t},\theta_{t}\}_{t\geq 0}$.
2:Initialization: ${\mathbf{x}}^{(0)}={\mathbf{z}}^{(0)}\in\mathcal{X}$, $\boldsymbol{\lambda}^{(0)}={\mathbf{0}}$.
3:for iteration $t=0,1,\dots,T-1$ do
4:  $\lambda_{i}^{(t+1)}=[\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)})]_{+},\;\forall\,i\in[m]$.
5:  ${\mathbf{x}}^{(t+1)}\approx\tilde{{\mathbf{x}}}^{(t+1)}:=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ that satisfies (23).
6:  ${\mathbf{z}}^{(t+1)}={\mathbf{z}}^{(t)}+\theta_{t}\cdot({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})$.
7:end for
At iteration $t$ of the iMELa method, the complexity for the APG method by Nesterov [37, Eq. (2.2.63)] to ensure (23) increases with the smoothness parameter of $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, which is $L+L\|\boldsymbol{\lambda}^{(t+1)}\|_{1}+p$. Fortunately, the lemma below shows that a uniform upper bound on $\{\|\boldsymbol{\lambda}^{(t)}\|\}_{t\geq 0}$ is available, whose proof is given in Section B.2.
Lemma 4.1.
The sequence $\{\boldsymbol{\lambda}^{(t)}\}_{t\geq 0}$ generated by Algorithm 1 satisfies, for any $t\geq 0$,
$\|\boldsymbol{\lambda}^{(t)}\|\leq M_{\boldsymbol{\lambda}}:=\max\left\{2\overline{\tau}B_{g},\frac{2\overline{\tau}(C_{\boldsymbol{\lambda}}+\overline{\tau}B_{g}^{2})}{2\underline{\tau}\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\textup{feas}})]}\right\},$ (24)
where $\overline{\tau}=\sup_{t\geq 0}\tau_{t}$, $\underline{\tau}=\inf_{t\geq 0}\tau_{t}$, and $C_{\boldsymbol{\lambda}}:=(B_{f}+pD_{\mathcal{X}}+1)\cdot D_{\mathcal{X}}$.
Thanks to Lemma 4.1 and the fact that $\|\boldsymbol{\lambda}^{(t+1)}\|_{1}\leq\sqrt{m}\|\boldsymbol{\lambda}^{(t+1)}\|\leq\sqrt{m}M_{\boldsymbol{\lambda}}$, function $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ is (pL)(p-L)-strongly convex and $K$-smooth in $\mathbf{x}$} over $\mathcal{X}$ for any $t\geq 0$, where
$K:=$ $\penalty 10000\ L+L\sqrt{m}M_{\boldsymbol{\lambda}}+p\geq L+L\|\boldsymbol{\lambda}^{(t+1)}\|_{1}+p.$ (25)
Given this result, the complexity for the APG method by Nesterov [37, Eq. (2.2.63)] to ensure (23) is $O(\ln(\epsilon_{t}^{-1}))$. This is stated in the following lemma whose proof is given in Section B.3.
Lemma 4.2.
Suppose the APG method by Nesterov [37, Eq. (2.2.63)] is applied to the subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ for $k_{t}$ iterations with any initial solution in $\mathcal{X}$ and returns a solution ${\mathbf{u}}^{(k_{t})}$. Let
${\mathbf{x}}^{(t+1)}=\textup{proj}_{\mathcal{X}}\big({\mathbf{u}}^{(k_{t})}-K^{-1}\cdot\nabla\mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\big).$
Then ${\mathbf{x}}^{(t+1)}$ satisfies (23) with
$k_{t}=O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right).$
In other words, (23) is ensured with complexity $O(\ln(\epsilon_{t}^{-1}))$ for any $t\geq 0$.
The procedure described in Lemma 4.2 is formally stated in Algorithm 2 in the appendix and is implemented in our numerical experiments.
5 Convergence Analysis for the iMELa Method
For simplicity of notation, let
$\sigma=(p-L)/p.$
The following lemma from Zhang and Luo [53, Lemma 3.5] characterizes the Lipschitz continuity of ${\mathbf{x}}({\mathbf{z}})$ in (17) and ${\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})$ in (20) under Assumption 3.1.
Lemma 5.1.
For any ${\mathbf{z}},{\mathbf{z}}^{\prime}\in\mathcal{X}$ and $\boldsymbol{\lambda}\geq{\mathbf{0}}$, we have
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\penalty 10000\ \sigma\|{\mathbf{x}}({\mathbf{z}})-{\mathbf{x}}({\mathbf{z}}^{\prime})\|,$ (26)
$\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|\geq$ $\penalty 10000\ \sigma\|{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}})-{\mathbf{x}}(\boldsymbol{\lambda},{\mathbf{z}}^{\prime})\|.$ (27)
Similar to the analysis in Zhang and Luo [53, 54], Zhang et al. [55], consider the potential function
$\hskip-7.22743pt\phi^{t}:=\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-2d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})+2v({\mathbf{z}}^{(t)}).$ (28)
By Assumption 3.1B, (16), (20) and (22), we have
$\phi^{t}=$ $\penalty 10000\ (\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))$
$+(v({\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}))+v({\mathbf{z}}^{(t)})$
$\geq$ $\penalty 10000\ v({\mathbf{z}}^{(t)})\geq\underline{f}.$ (29)
Following the analysis in Zhang and Luo [53, 54], Zhang et al. [55], we characterize the change of $\phi^{t}$ after each iteration as follows. The proof is given in Section B.4.
Proposition 5.2.
Suppose that in Algorithm 1,
$\tau_{t}=\tau=\frac{p-L}{4B_{g}^{2}}\text{ and }\theta_{t}=\theta\leq\frac{p-L}{18p},\penalty 10000\ \forall\,t\geq 0.$
The sequence $\{({\mathbf{x}}^{(t)},\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$\phi^{t}-\phi^{t+1}\geq$ $\penalty 10000\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$+\frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\,\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}.$ (30)
According to Proposition 5.2, to ensure a sufficient decrease of $\phi_{t}$, it is necessary to control the negative term $6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ appearing in (30). To this end, we employ the local EBC with exponent $d$ in Assumption 3.3. Specifically, we show in the following proposition that, when $\tilde{{\mathbf{x}}}^{(t+1)}$ is a near-KKT point, the local EBC allows us to upper bound $\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ using a multiple of $\text{dist}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))^{2/d}$. The proof is presented in Section C.1.
Proposition 5.3.
Suppose that Assumption 3.3 holds. There exists a constant $R(\delta)$, depending only on the constant $\delta$ in Assumption 3.3, such that, under the same conditions as in Proposition 5.2, if
$\textup{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big)^{2}$
$\quad+\penalty 10000\ \textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}\leq R^{2}(\delta),$ (31)
it holds that
$\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\penalty 10000\ \frac{\gamma^{1/d}K}{p-L}\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2/d}.$ (32)
Thanks to Proposition 5.3, when the local regularity condition (31) holds, the negative term $6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ in (30) can be bounded above by a multiple of the quantity $\text{dist}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))^{2/d}$ as long as $\theta_{t}=\theta$ is chosen sufficiently small. This bound follows directly from the local EBC with exponent $d$ in Assumption 3.3.
However, when (31) does not hold, the local EBC is unavailable. In this case, we resort to the following global result, which holds without any error bound assumption. See Section C.2 for its proof.
Lemma 5.4.
With $M_{\boldsymbol{\lambda}}$ defined in (24), it holds that
$\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq\frac{2M_{\boldsymbol{\lambda}}}{p-L}\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big).$ (33)
The result of Lemma 5.4 can be viewed as a global analogue of Assumption 3.3 with exponent $d=2$, and serves as a fallback estimate when condition (31) fails. Using Proposition 5.3 and Lemma 5.4, we can ensure a sufficient decrease of $\phi_{t}$ by choosing $\theta_{t}=\theta$ and $\tau_{t}$ in Algorithm 1 carefully. This result is presented below and is the key to establish the total complexity. The proof is given in Section C.3.
Proposition 5.5.
Suppose that Assumptions 3.1 and 3.3 hold and, in Algorithm 1, we set
$\begin{aligned} \tau_{t}=\tau=&\penalty 10000\ \frac{p-L}{4B_{g}^{2}},\\ \theta_{t}=\theta\leq&\penalty 10000\ \tilde{\theta}_{d}:=\min\left\{\frac{p-L}{18p},6p\tau,\frac{R^{2}(\delta)(p-L)^{2}\tau}{2\cdot 12^{3}p^{3}M_{\boldsymbol{\lambda}}^{2}},\hat{\theta}_{d}\right\},\end{aligned}$ (34)
for $\forall\,t\geq 0$, where $C_{d}:=\frac{\gamma^{1/d}K}{p-L}$ and
$\hat{\theta}_{d}:=\frac{\tau}{12pC_{d}}\left(\frac{\tau}{12p}\right)^{1-1/d}.$ (35)
Then the sequence $\{({\mathbf{x}}^{(t)},\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$\phi^{t}-\phi^{t+1}\geq$ $\penalty 10000\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})\|^{2}+\frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$+\frac{\tau}{2}\textup{dist}\big({\mathbf{g}}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\Xi_{d}(\theta)-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}},$ (36)
where
$\Xi_{d}(\theta):=\begin{cases}\left(1-\dfrac{1}{d}\right)\left(\dfrac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta\,C_{d}\big)^{\frac{d}{d-1}},&d\in(1,2],\\[4.30554pt] 0,&d=1.\end{cases}$ (37)
By Proposition 5.5, we establish the complexity of Algorithm 1 in the following theorem whose proof is in Section D.
Theorem 5.6.
Suppose that the assumptions in Proposition 5.5 hold with a fixed exponent $d\in[1,2]$. For a given target accuracy $\epsilon>0$, choose $\theta_{t}=\theta=\theta(\epsilon)$, where
$\theta(\epsilon):=\min\left\{\tilde{\theta}_{d},\,\frac{6p\tau}{M_{\boldsymbol{\lambda}}^{2}+1},\,\left(\frac{\epsilon^{2}}{2A_{d}}\right)^{d-1}\right\},\penalty 10000\ \text{ with }\penalty 10000\ A_{d}:=24p\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}(6pC_{d})^{\frac{d}{d-1}},$
and $\tilde{\theta}_{d}$, $C_{d}$ (and $\tau$) defined in Proposition 5.5. If $\epsilon_{t}=\frac{c}{t+1}$ for any constant $c>0$ and
$\begin{aligned} T:=\left\lceil\frac{2\mathcal{C}_{0}(\theta(\epsilon))}{\epsilon^{2}}\right\rceil,\text{ where }\penalty 10000\ \mathcal{C}_{0}(\theta):=&\penalty 10000\ \left(\begin{array}[]{l}\left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}\\ +\;\mathcal{C}_{1}(\theta)\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}\right)\end{array}\right),\\ \mathcal{C}_{1}(\theta):=&\penalty 10000\ \max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\},\end{aligned}$ (38)
Algorithm 1 finds an $\epsilon$-KKT solution for (1) with oracle complexity $O(\epsilon^{-2d}\ln(\epsilon^{-1}))$.
6 Numerical Experiment
We demonstrate the performance of the iMELa method on a fairness-aware classification problem, which is an instance of (1). We compare it with the inexact proximal point penalty (iPPP) method [28], the damped proximal augmented Lagrangian method (DPALM) [8],the smoothed proximal Lagrangian method (SP-LM) [40], and the switching subgradient (SSG) method [16].
First, we define the instance of (1) solved in our experiments. Given a feature vector ${\mathbf{a}}\in\mathbb{R}^{d}$ and a class label $b\in\{1,-1\}$, the goal of linear binary classification is to learn a model ${\mathbf{x}}\in\mathbb{R}^{d}$ to predict $b$ based on the score ${\mathbf{x}}^{\top}{\mathbf{a}}$ with a larger score indicating a higher chance of a positive label. Suppose that a training set $\mathcal{D}=\{({\mathbf{a}}_{i},b_{i})\}_{i=1}^{n}$ is available. A model $\mathbf{x}$} can be obtained by solving
$\mathcal{L}^{*}:=\min_{{\mathbf{x}}\in\mathcal{X}}\Big\{\mathcal{L}({\mathbf{x}}):=\frac{1}{n}\sum_{i=1}^{n}\ell(b_{i}\cdot{\mathbf{x}}^{\top}{\mathbf{a}}_{i})\Big\},$ (39)
where $\ell(z)=\log(1+\exp(-z))$ and $\mathcal{X}$ is a convex and compact set.
However, solving (39) only optimizes the classification accuracy of the resulting model but does not guarantee its fairness. Suppose that a data point $\mathbf{a}$} is classified as positive if ${\mathbf{x}}^{\top}{\mathbf{a}}\geq 0$ and as negative otherwise. Also, suppose that model $\mathbf{x}$} is applied to two groups of data, a protected group $\mathcal{D}_{p}=\{{\mathbf{a}}_{i}^{p}\}_{i=1}^{n_{p}}$ and an unprotected group $\mathcal{D}_{u}=\{{\mathbf{a}}_{i}^{u}\}_{i=1}^{n_{u}}$. A measure of the fairness of $\mathbf{x}$} between $\mathcal{D}_{p}$ and $\mathcal{D}_{u}$ is defined as
$\Big|\frac{1}{n_{p}}\sum_{i=1}^{n_{p}}\mathbb{I}({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{p}\geq 0)-\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}\mathbb{I}({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{u}\geq 0)\Big|,$
where $\mathbb{I}(\cdot)$ is the one-zero indicator function. This measure is known as the demographic parity [9]. When its value is small, the model $\mathbf{x}$} produces similar predicted positive rates in both groups, indicating the fairness of the model. However, this measure is computationally challenging because of the discontinuity of $\mathbb{I}(\cdot)$. Hence, we approximate this measure by a continuous function of $\mathbf{x}$} defined as
$\mathcal{R}({\mathbf{x}}):=\frac{1}{n_{p}}\sum_{i=1}^{n_{p}}\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{p})-\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{u}),$ (40)
where $\sigma(z)=\exp(z)/(1+\exp(z))$.
a9a bank COMPAS
Objective Refer to caption Refer to caption Refer to caption
Infeasibility Refer to caption Refer to caption Refer to caption
Stationarity Refer to caption Refer to caption Refer to caption
Complementary Slackness Refer to caption Refer to caption Refer to caption
Figure 1: Performances of iMELa, SP-LM, iPPP, DPALM, and SSG with static and diminishing step-sizes vs. number of steps on classification problems with demographic parity fairness.
To obtain a fair $\mathbf{x}$}, we balance the loss $\mathcal{L}({\mathbf{x}})$ in (39) and the fairness measure $|\mathcal{R}({\mathbf{x}})|$ in (40) by solving
$\min_{{\mathbf{x}}\in\mathcal{X}}\frac{1}{2}\big(\mathcal{R}({\mathbf{x}})\big)^{2}\text{ s.t. }\mathcal{L}({\mathbf{x}})\leq\mathcal{L}^{*}+\kappa,$ (41)
where $\kappa$ is the slackness parameter indicating how much we are willing to increase the loss in order to reduce $|\mathcal{R}({\mathbf{x}})|$ to obtain a more fair model, and $\mathcal{X}$ is the same as that in (39). In the experiments, we set $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:\|{\mathbf{x}}\|_{1}\leq r\}$ with $r=6\|{\mathbf{a}}_{i}\|_{1}$ to satisfy Assumption 3.1A.
We now briefly justify that problem (41) satisfies Assumption 3.3 with exponent $d=1$. Specifically, we show that the constraint system satisfies the local LICQ condition at all KKT points, which implies the desired local EBC with exponent $d=1$ via Proposition 3.4. The formal statement and proof are deferred to Section E.2.
We solve problem (41) on three datasets: a9a [18], bank [35] and COMPAS [1]. Details about these datasets and how they are used to formulate (41) are provided in Section E.1. The implementation details of each method, including the choices of tuning parameters, are provided in Section E.3.
We compare the methods based on the objective value, infeasibility, stationarity, and complementary slackness (see (5), (6) and (7)) achieved by ${\mathbf{x}}^{(t)}$ and the corresponding Lagrangian multipliers. [^3] Since the SSG method does not generate a Lagrangian multiplier, we do not report its measures of stationarity and complementary slackness. While the calculations for most of these metrics are straightforward, additional details are provided for computing the stationarity metric. Observe that $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:{\mathbf{B}}{\mathbf{x}}\leq r\cdot\mathbf{1}\}$, where ${\mathbf{B}}\in\mathbb{R}^{2^{d}\times d}$ is a matrix whose rows are all vectors in $\{1,-1\}^{d}$, and $\mathbf{1}$ denotes the all-one vector in $\mathbb{R}^{2^{d}}$. If $\|{\mathbf{x}}\|_{1}<r$, then $\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=\{{\mathbf{0}}\}$ so that the stationarity measure in (5) reduces to $\|\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})\|$, where $\lambda\geq 0$ is the Lagrange multiplier produced by the algorithm to pair with $\mathbf{x}$}. If $\|{\mathbf{x}}\|_{1}=r$, then by Hiriart-Urruty and Lemaréchal [13, Example A.5.2.6 (b)], there exists an index set of the active constraints at ${\mathbf{x}}\in\mathcal{X}$ defined by
$J({\mathbf{x}})=\{j=1,\dots,2^{d}:[{\mathbf{B}}]_{j}{\mathbf{x}}=r\}$ (42)
where ${\mathbf{B}}]_{j}$ denotes the $j$-th row of $\mathbf{B}$}. Then the normal cone to $\mathcal{X}$ at $\mathbf{x}$} in this case is
$\mathcal{N}_{\mathcal{X}}({\mathbf{x}})=$ $\penalty 10000\ \text{cone}\{[{\mathbf{B}}]_{j}^{\top}:j\in J({\mathbf{x}})\}$
$=$ $\penalty 10000\ \textstyle\{\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}:a_{j}\geq 0,\,j\in J({\mathbf{x}})\}.$ (43)
Therefore, the stationarity measure in (5) can be calculated by solving
$\min_{a_{j}\geq 0,\,j\in J({\mathbf{x}})}\left\|\mathcal{R}({\mathbf{x}})\cdot\nabla\mathcal{R}({\mathbf{x}})+\lambda\cdot\nabla\mathcal{L}({\mathbf{x}})+\textstyle\sum_{j\in J({\mathbf{x}})}a_{j}[{\mathbf{B}}_{j}]^{\top}\right\|.$
We use mpcActiveSetSolver in MATLAB to solve this quadratic program.
To compare the methods in their oracle complexity, in Figure 1, we plot the four aforementioned performance metrics against the accumulated number of gradient steps performed by the APG method called within the iMELa method, the iPPP method, and DPALM, and the number of gradient steps performed by SP-LM and the SSG method. According to Figure 1, on all the instances studied, the iMELa method performs similarly to SP-LM and outperforms the iPPP method in all performance metrics. Moreover, DPALM exhibits performance comparable to iMELa and SP-LM on COMPAS, while on a9a and bank it converges more slowly and attains worse infeasibility metrics. These observations are consistent with our theoretical finding that the iMELa method has an oracle complexity similar to SP-LM and better than those of the iPPP method and DPALM. In addition, the iMELa method and SP-LM reduce the infeasibility quickly and maintain a (nearly) feasible solution throughout the remaining iterations. Finally, we find that for all instances, the number of gradient steps performed by the APG method for solving (116) within the iMELa method and for solving (118) within DPALM is generally much smaller than that for solving (117), which is the main reason for the better empirical performance of the iMELa method and DPALM than the iPPP method.
7 Conclusion
We analyze the oracle complexity of the iMELa method for finding an $\epsilon$-KKT point in smooth non-convex functional constrained optimization. Each iteration consists of approximately solving a strongly convex subproblem, followed by updates to the Lagrange multipliers and the proximal center. Our results show that the oracle complexity of the iMELa method matches the best known result in the literature, up to a logarithmic factor, under a local error bound condition with exponent $d=1$, which is a strictly weaker assumption than the local LICQ condition required to achieve the best complexity guarantee.
Appendix A Proof of Proposition 3.4
In this section, we provide the detailed proof for Proposition 3.4, starting with a few technical lemmas.
Lemma A.1.
The set $\mathcal{X}^{*}$ is compact.
Proof.
By the structure of $\mathcal{X}$ and the Slater’s condition in Assumption 3.1A and F, $\mathcal{X}^{*}$ is non-empty (see, e.g. Bertsekas [2, Sec. 3.3]). We then prove the lemma by two steps.
First, we show that, for any KKT point ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, its any associated vector of Lagrangian multipliers $\boldsymbol{\lambda}^{*}$ satisfies that $\|\boldsymbol{\lambda}^{*}\|\leq B_{\boldsymbol{\lambda}}:=\frac{B_{f}D_{\mathcal{X}}}{\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]}$. By the KKT conditions at $({\mathbf{x}}^{*},\boldsymbol{\lambda}^{*})$ in (1), we have
$\textstyle\boldsymbol{\xi}^{*}:=-\nabla f({\mathbf{x}}^{*})-\sum_{i=1}^{m}\lambda_{i}^{*}\nabla g_{i}({\mathbf{x}}^{*})\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*})\text{ and }\lambda_{i}^{*}g_{i}({\mathbf{x}}^{*})=0,\;i\in[m].$ (44)
Hence, by the convexity of $g_{i}$’s and $\boldsymbol{\lambda}^{*}\geq{\mathbf{0}}$, it follows that
$\|\boldsymbol{\lambda}^{*}\|_{1}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]\leq$ $-\textstyle\sum_{i=1}^{m}\lambda_{i}^{*}g_{i}({\mathbf{x}}_{\text{feas}})\leq-\textstyle\sum_{i=1}^{m}\lambda_{i}^{*}\left(g_{i}({\mathbf{x}}^{*})+\langle\nabla g_{i}({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\rangle\right)$
$=$ $\penalty 10000\ \big\langle\boldsymbol{\xi}^{*}+\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big\rangle\leq\big\langle\nabla f({\mathbf{x}}^{*}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{*}\big\rangle\leq B_{f}D_{\mathcal{X}},$
where ${\mathbf{x}}_{\text{feas}}$ is given in Assumption 3.1F, the equality holds by (44), the third inequality holds by $\boldsymbol{\xi}^{*}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{*})$, and the last one is from Assumption 3.1A, D, and F. Thus, we obtain $\|\boldsymbol{\lambda}^{*}\|\leq\|\boldsymbol{\lambda}^{*}\|_{1}\leq B_{\boldsymbol{\lambda}}$, which proves the claimed result.
Second, by Assumption 3.1A, $\mathcal{X}^{*}$ is bounded. Hence, it suffices to show that $\mathcal{X}^{*}$ is closed. Suppose there is a sequence of KKT points and their associated vector of Lagrangian multipliers, denote by $\{(\bar{{\mathbf{x}}}^{(k)},\bar{\boldsymbol{\lambda}}^{(k)})\}_{k\geq 1}\subset\mathcal{X}^{*}\times\mathbb{R}_{+}^{m}$, such that $\bar{{\mathbf{x}}}^{(k)}\rightarrow\bar{\mathbf{x}}^{*}$ as $k\rightarrow\infty$ for some $\bar{\mathbf{x}}^{*}\in\mathcal{X}$. By definition, there exists $\bar{\boldsymbol{\xi}}^{(k)}\in\mathcal{N}_{\mathcal{X}}(\bar{{\mathbf{x}}}^{(k)})$ such that
$\nabla f(\bar{{\mathbf{x}}}^{(k)})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{(k)}\nabla g_{i}(\bar{{\mathbf{x}}}^{(k)})+\bar{\boldsymbol{\xi}}^{(k)}={\mathbf{0}},\penalty 10000\ g_{i}(\bar{{\mathbf{x}}}^{(k)})\leq 0,\penalty 10000\ \bar{\lambda}_{i}^{(k)}g_{i}(\bar{{\mathbf{x}}}^{(k)})=0,\penalty 10000\ \forall\,i\in[m].$
Since $\|\bar{\boldsymbol{\lambda}}^{(k)}\|\leq B_{\boldsymbol{\lambda}}$, passing to a subsequence if necessary, it holds that $\lim_{k\rightarrow\infty}\bar{\boldsymbol{\lambda}}^{(k)}=\bar{\boldsymbol{\lambda}}^{*}$ for some $\bar{\boldsymbol{\lambda}}^{*}\geq{\mathbf{0}}$. As $k\rightarrow\infty$, we conclude from the continuity of $\nabla f(\cdot)$ and $\nabla{\mathbf{g}}(\cdot)$ and the outer semi-continuity of $\mathcal{N}_{\mathcal{X}}(\cdot)$ [46, Proposition 6.6] that $\bar{\boldsymbol{\xi}}^{(k)}\rightarrow\bar{\boldsymbol{\xi}}^{*}$ for some $\bar{\boldsymbol{\xi}}^{*}\in\mathcal{N}_{\mathcal{X}}(\bar{{\mathbf{x}}}^{*})$ and
$\nabla f(\bar{{\mathbf{x}}}^{*})+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}^{*}\nabla g_{i}(\bar{{\mathbf{x}}}^{*})+\bar{\boldsymbol{\xi}}^{*}={\mathbf{0}},\penalty 10000\ g_{i}(\bar{{\mathbf{x}}}^{*})\leq 0,\penalty 10000\ \bar{\lambda}_{i}^{*}g_{i}(\bar{{\mathbf{x}}}^{*})=0,\penalty 10000\ \forall\,i\in[m],$
meaning that $\bar{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$, which completes the proof. ∎
Lemma A.2.
For any $\vartheta>0$, there exists $r(\vartheta)>0$ such that, if $\mathbf{x}$} is feasible to (1) and $\textup{dist}({\mathbf{x}},\mathcal{X}^{*})\leq r(\vartheta)$, then there must exist ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\vartheta$ and $J_{g}({\mathbf{x}})\subset J_{g}({\mathbf{x}}^{*})\text{ and }J_{A}({\mathbf{x}})\subset J_{A}({\mathbf{x}}^{*})$.
Proof.
We prove this by contradiction. Suppose that the claim does not hold. Then there exists $\vartheta>0$ and a sequence $\{{\mathbf{x}}^{(k)}\}_{k\geq 1}$ such that ${\mathbf{x}}^{(k)}$ is feasible to (1) and $\text{dist}({\mathbf{x}}^{(k)},\mathcal{X}^{*})\leq r_{k}:=\frac{\vartheta}{k}\leq\vartheta$ but $J_{g}({\mathbf{x}}^{(k)})\not\subset J_{g}({\mathbf{x}}^{*})\text{ or }J_{A}({\mathbf{x}}^{(k)})\not\subset J_{A}({\mathbf{x}}^{*})$ for any $k$ and any ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ satisfying $\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{*}\|\leq\vartheta$.
Due to the compactness of $\mathcal{X}$ and $\mathcal{X}^{*}$ (see Lemma A.1), by passing to a subsequence if necessary, there exists some $\bar{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$ such that ${\mathbf{x}}^{(k)}\rightarrow\bar{{\mathbf{x}}}^{*}$ as $k\rightarrow\infty$. By the continuity of ${\mathbf{g}}(\cdot)$ and ${\mathbf{A}}(\cdot)-{\mathbf{b}}$, it holds that for any $i\in[m]$, if $g_{i}(\bar{{\mathbf{x}}}^{*})<0$, then $g_{i}({\mathbf{x}}^{(k)})<0$ as well when $k$ is large enough, and for any $j\in[l]$, if ${\mathbf{A}}\bar{{\mathbf{x}}}^{*}-{\mathbf{b}}]_{j}<0$, then ${\mathbf{A}}{\mathbf{x}}^{(k)}-{\mathbf{b}}]_{j}<0$ as well when $k$ is large enough. That means $J_{g}({\mathbf{x}}^{(k)})\subset J_{g}(\bar{\mathbf{x}}^{*})$ and $J_{A}({\mathbf{x}}^{(k)})\subset J_{A}(\bar{\mathbf{x}}^{*})$ for a large enough $k$. This contradicts to the assumption. Therefore, the conclusion holds. ∎
Lemma A.3.
Suppose the constraints of (1) satisfy the local LICQ in (15). Then there exists $\delta^{\prime}>0$ such that, if $\mathbf{x}$} is feasible to (1) and $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta^{\prime}$ for some ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, then $\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)\geq\frac{\zeta}{2}$ for any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subset J_{A}({\mathbf{x}}^{*})$.
Proof.
Fix ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ and any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$, $I_{A}\subset J_{A}({\mathbf{x}}^{*})$. By the continuity of $\sigma_{\min}(\cdot)$ (see e.g. Horn and Johnson [15, Corollary 7.3.5 (a)]), for any $\mathbf{x}$} we have
$\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]-\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big\|$
$=$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),\;{\mathbf{O}}\big]\big\|$
$\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-\big\|\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*})\big\|.$
Since the functions $g_{i}$, $i\in[m]$, are $L$-smooth by Assumption 3.1C, we have
$\big\|\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}})-\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*})\big\|\leq L\sqrt{|I_{g}|}\,\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq L\sqrt{m}\,\|{\mathbf{x}}-{\mathbf{x}}^{*}\|.$
Define $L_{H}:=L\sqrt{m}$ and set
$\delta^{\prime}:=\frac{\zeta}{2L_{H}}=\frac{\zeta}{2L\sqrt{m}}.$
Then for all feasible $\mathbf{x}$} satisfying $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta^{\prime}$, we obtain
$\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)\geq$ $\penalty 10000\ \sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{I_{g}}({\mathbf{x}}^{*}),{\mathbf{A}}_{I_{A}}^{\top}\big]\big)-L_{H}\|{\mathbf{x}}-{\mathbf{x}}^{*}\|$
$\geq$ $\penalty 10000\ \zeta-L_{H}\delta^{\prime}=\zeta-\tfrac{\zeta}{2}=\tfrac{\zeta}{2}.$
Proof of Proposition 3.4.
Let $\delta=\min\left\{\frac{r(\delta^{\prime})}{2},\frac{\zeta}{2L\sqrt{m}}\right\}$, where $r(\cdot)$ is from Lemma A.2 and $\delta^{\prime}$ is from Lemma A.3. Suppose that ${\mathbf{x}}\in\mathcal{X}$ and $\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta$ for some ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$, we then need to show that there exists $\gamma>0$ such that (14) holds with $d=1$ for any $I_{g}\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}\subset J_{A}({\mathbf{x}}^{*})$. For such $I_{g}$ and $I_{A}$, consider the minimization problem
$\bar{{\mathbf{x}}}=\mathop{\mathrm{arg\,min}{}}_{{\mathbf{z}}}\frac{1}{2}\|{\mathbf{z}}-{\mathbf{x}}\|^{2}\text{ s.t.}$ $\begin{array}[]{l}g_{i}({\mathbf{z}})=0,\,i\in I_{g},\,g_{i}({\mathbf{z}})\leq 0,\,i\in[m]\backslash I_{g},\\[0.43057pt] [{\mathbf{A}}{\mathbf{z}}-{\mathbf{b}}]_{j}=0,\,j\in I_{A},\,[{\mathbf{A}}{\mathbf{z}}-{\mathbf{b}}]_{j}\leq 0,\,j\in[l]\backslash I_{A}.\end{array}$ (47)
Note that $\bar{{\mathbf{x}}}$ is well-defined because the feasible set of (47) is compact and non-empty as it at least contains $\mathbf{x}$$^{*}$. As a result, $\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},I_{A}))=\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|\leq\|{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta$ and thus $\text{dist}(\bar{{\mathbf{x}}},\mathcal{X}^{*})\leq\|\bar{{\mathbf{x}}}-{\mathbf{x}}^{*}\|\leq 2\delta\leq r(\delta^{\prime})$. Since $\bar{\mathbf{x}}$ is also feasible to (1), by invoking Lemma A.2, there exists $\bar{\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|\bar{\mathbf{x}}-\bar{\mathbf{x}}^{*}\|\leq\delta^{\prime}$ and $J_{g}(\bar{\mathbf{x}})\subset J_{g}(\bar{\mathbf{x}}^{*})$ and $J_{A}(\bar{\mathbf{x}})\subset J_{A}(\bar{\mathbf{x}}^{*})$. By Lemma A.3 at $\bar{\mathbf{x}}$ and $\delta^{\prime}$, we have $\sigma_{\min}\big(\big[\nabla{\mathbf{g}}_{J_{g}(\bar{\mathbf{x}})}(\bar{\mathbf{x}}),{\mathbf{A}}_{J_{A}(\bar{\mathbf{x}})}^{\top}\big]\big)\geq\frac{\zeta}{2}$. This means (47) satisfies the LICQ at $\bar{\mathbf{x}}$. Hence, there exist a vector of Lagrangian multipliers $\bar{\boldsymbol{\lambda}}\in\mathbb{R}^{m}$ and $\bar{\mathbf{w}}\in\mathbb{R}^{l}$ such that
$\bar{{\mathbf{x}}}-{\mathbf{x}}+\textstyle\sum_{i=1}^{m}\bar{\lambda}_{i}\cdot\nabla g_{i}(\bar{{\mathbf{x}}})+\textstyle\sum_{j=1}^{l}\bar{w}_{j}\cdot[{\mathbf{A}}]_{j}^{\top}={\mathbf{0}},$ (48)
$\bar{\lambda}_{i}\geq 0,\;i\in[m]\backslash I_{g},\quad\bar{w}_{j}\geq 0,\;j\in[l]\backslash I_{A},$ (49)
$\bar{\lambda}_{i}\cdot g_{i}(\bar{{\mathbf{x}}})=0,\;i\in[m]\backslash I_{g},\quad\bar{w}_{j}\cdot[{\mathbf{A}}\bar{{\mathbf{x}}}-{\mathbf{b}}]_{j}=0,\;j\in[l]\backslash I_{A}.$ (50)
Notice $I_{g}\subset J_{g}(\bar{{\mathbf{x}}})$ and $I_{A}\subset J_{A}(\bar{{\mathbf{x}}})$. By (50), we must have $\bar{\lambda}_{i}=0$ for $i\in[m]\backslash J_{g}(\bar{{\mathbf{x}}})$ and $\bar{w}_{j}=0$ for $j\in[l]\backslash J_{A}(\bar{{\mathbf{x}}})$. Thus, (48) and $\sigma_{\min}([\nabla{\mathbf{g}}_{J_{g}(\bar{{\mathbf{x}}})}(\bar{{\mathbf{x}}}),{\mathbf{A}}_{J_{A}(\bar{{\mathbf{x}}})}^{\top}])\geq\textstyle\frac{\zeta}{2}$ imply
[…]
Moreover, for $i\in I_{g}$, by the $L$-smoothness of $g_{i}$ and the fact that $g_{i}(\bar{{\mathbf{x}}})=0$, we have
$\bar{\lambda}_{i}\cdot\left\langle\nabla g_{i}(\bar{{\mathbf{x}}}),{\mathbf{x}}-\bar{{\mathbf{x}}}\right\rangle\leq$ $\penalty 10000\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})-g_{i}(\bar{{\mathbf{x}}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}$
$=$ $\penalty 10000\ |\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})|+\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}.$ (52)
[…]
$\bar{\lambda}_{i}\cdot\left\langle\nabla g_{i}(\bar{{\mathbf{x}}}),{\mathbf{x}}-\bar{{\mathbf{x}}}\right\rangle\leq\bar{\lambda}_{i}\cdot\left(g_{i}({\mathbf{x}})-g_{i}(\bar{{\mathbf{x}}})\right)\leq|\bar{\lambda}_{i}|\cdot[g_{i}({\mathbf{x}})]_{+}.$ (53)
Following similar arguments, we have
[…]
Taking the inner product of the left-hand side of (48) and $\bar{{\mathbf{x}}}-{\mathbf{x}}$ and applying (52), (53), and (54) lead to
$\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|^{2}\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|\bar{\lambda}_{i}|\cdot|g_{i}({\mathbf{x}})|+\sum_{i\in[m]\backslash I_{g}}|\bar{\lambda}_{i}|\cdot[g_{i}({\mathbf{x}})]_{+}\\[3.87498pt] +\sum_{j\in I_{A}}|\bar{w}_{j}|\cdot\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|+\sum_{j\in[l]\backslash I_{A}}|\bar{w}_{j}|\cdot\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\end{array}\right)+\sum_{i\in I_{g}}\frac{|\bar{\lambda}_{i}|L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}$
$\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\[1.93748pt] +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\sqrt{\|\bar{\boldsymbol{\lambda}}\|^{2}+\|\bar{{\mathbf{w}}}\|^{2}}$
$+\sqrt{m}\sqrt{\|\bar{\boldsymbol{\lambda}}\|^{2}+\|\bar{{\mathbf{w}}}\|^{2}}\cdot\frac{L}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}.$
Recall (51) and the fact that $\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|\leq\delta\leq\frac{\zeta}{2L\sqrt{m}}$. The inequality above implies
$\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|^{2}\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\[1.93748pt] +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\frac{2\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|}{\zeta}$
$+\sqrt{m}\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|\cdot\frac{L}{\zeta}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2}$
$\leq$ $\left(\begin{array}[]{l}\sum_{i\in I_{g}}|g_{i}({\mathbf{x}})|^{2}+\sum_{i\in[m]\backslash I_{g}}\left([g_{i}({\mathbf{x}})]_{+}\right)^{2}\\[1.93748pt] +\sum_{j\in I_{A}}\left|[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right|^{2}+\sum_{j\in[l]\backslash I_{A}}\left(\left[[{\mathbf{A}}{\mathbf{x}}-{\mathbf{b}}]_{j}\right]_{+}\right)^{2}\end{array}\right)^{1/2}\cdot\frac{2\|\bar{{\mathbf{x}}}-{\mathbf{x}}\|}{\zeta}+\frac{1}{2}\|{\mathbf{x}}-\bar{{\mathbf{x}}}\|^{2},$
which implies (14) with $d=1$ and $\gamma=16/\zeta^{2}$.
On the contrary, we can next show that Assumption 3.3 does not imply the local LICQ in (15), indicating that Assumption 3.3 is weaker than the local LICQ.
Consider an instance with an objective $f({\mathbf{x}})=x_{1}^{2}+x_{2}^{2}$ and $\mathcal{X}:=\{{\mathbf{x}}\in\mathbb{R}^{2}:\ -2\leq x_{1}\leq 2,\ -2\leq x_{2}\leq 2\}$, and introduce two nonlinear convex inequality constraints $g_{1}({\mathbf{x}}):=x_{1}$, $g_{2}({\mathbf{x}}):=x_{1}+x_{1}^{2}$. Clearly, both $g_{1}$ and $g_{2}$ are smooth and convex on $\mathbb{R}^{2}$, and the system admits a Slater point, e.g., ${\mathbf{x}}_{\text{feas}}:=(-1/2,0)^{\top}$ satisfying $g_{1}({\mathbf{x}}_{\text{feas}})<0$ and $g_{2}({\mathbf{x}}_{\text{feas}})<0$. It is immediate that ${\mathbf{x}}^{*}={\mathbf{0}}$ is the unique KKT point of this instance. Moreover, we have $J_{g}({\mathbf{x}}^{*})=\{1,2\}$ and $J_{A}({\mathbf{x}}^{*})=\emptyset$. However, the local LICQ fails at $\mathbf{x}$$^{*}$. Indeed, both constraints are active at $\mathbf{x}$$^{*}$, while their gradients are linearly dependent:
$\nabla g_{1}({\mathbf{x}}^{*})=(1,0)^{\top},\quad\nabla g_{2}({\mathbf{x}}^{*})=(1+2x_{1},0)^{\top}\big|_{{\mathbf{x}}={\mathbf{x}}^{*}}=(1,0)^{\top},$
so the active gradients cannot be linearly independent.
On the other hand, (14) holds around $\mathbf{x}$$^{*}$ with exponent $d=1$ and an explicit constant $\gamma=4$. Fix $\delta:=1/2$. For any ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$, we have $x_{1}\in[-1/2,1/2]$. Now take any $I_{g}\subseteq J_{g}({\mathbf{x}}^{*})=\{1,2\}$ and $I_{A}\subseteq J_{A}({\mathbf{x}}^{*})=\emptyset$.
Case 1: $I_{g}=\emptyset$. Then, by definition, $\mathcal{S}(\emptyset,\emptyset)=\{{\mathbf{x}}\in\mathcal{X}:\ g_{1}({\mathbf{x}})\leq 0,\ g_{2}({\mathbf{x}})\leq 0\}$. For any ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$, if $x_{1}\leq 0$ then $x_{1}\in[-1/2,0]$ and hence $1+x_{1}\geq 1/2$, which yields $g_{2}({\mathbf{x}})=x_{1}+x_{1}^{2}=x_{1}(1+x_{1})\leq 0$. Hence, for any ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$, we have $g_{1}({\mathbf{x}})\leq 0\Rightarrow g_{2}({\mathbf{x}})\leq 0$, and thus $\{{\mathbf{x}}\in\mathcal{X}:\ x_{1}\leq 0\}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)\subseteq\mathcal{S}(\emptyset,\emptyset)$. Conversely, ${\mathbf{x}}\in\mathcal{S}(\emptyset,\emptyset)$ implies $g_{1}({\mathbf{x}})\leq 0$, i.e., $x_{1}\leq 0$. Therefore, for all ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$,
$\text{dist}({\mathbf{x}},\mathcal{S}(\emptyset,\emptyset))^{2}=([g_{1}({\mathbf{x}})]_{+})^{2}\leq 4\big(([g_{1}({\mathbf{x}})]_{+})^{2}+([g_{2}({\mathbf{x}})]_{+})^{2}\big).$
Case 2: $I_{g}\neq\emptyset$. We show that for all ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$, $\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))=|x_{1}|$.
If $1\in I_{g}$, then every ${\mathbf{z}}\in\mathcal{S}(I_{g},\emptyset)$ satisfies $g_{1}({\mathbf{z}})=0$, and hence $z_{1}=0$. Therefore,
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))\geq\text{dist}({\mathbf{x}},\{{\mathbf{z}}\in\mathcal{X}:\ z_{1}=0\})=|x_{1}|.$
On the other hand, the point $\hat{{\mathbf{x}}}:=(0,x_{2})^{\top}$ belongs to $\mathcal{S}(I_{g},\emptyset)$, which yields
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))\leq\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|=|x_{1}|.$
Combining the two inequalities gives $\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))=|x_{1}|$.
If $1\notin I_{g}$, then $2\in I_{g}$ and any ${\mathbf{z}}\in\mathcal{S}(I_{g},\emptyset)$ must satisfy $g_{2}({\mathbf{z}})=0$, i.e., $z_{1}\in\{0,-1\}$. Since $x_{1}\in[-1/2,1/2]$ on $\mathbb{B}({\mathbf{x}}^{*},\delta)$, we have $|x_{1}-z_{1}|\geq|x_{1}|$ for all $z_{1}\in\{0,-1\}$, and hence
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))\geq|x_{1}|.$
Moreover, the point $\hat{{\mathbf{x}}}:=(0,x_{2})^{\top}$ satisfies $g_{2}(\hat{{\mathbf{x}}})=0$ and $g_{1}(\hat{{\mathbf{x}}})=0$ and thus $\hat{{\mathbf{x}}}\in\mathcal{S}(I_{g},\emptyset)$, which implies
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))\leq\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|=|x_{1}|.$
Combining the upper and lower bounds yields $\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))=|x_{1}|$.
With the distance identity established, we verify (14) with $\gamma=4$. If $x_{1}\geq 0$, then $g_{1}({\mathbf{x}})]_{+}=x_{1}$ and thus
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))^{2}=|x_{1}|^{2}=([g_{1}({\mathbf{x}})]_{+})^{2}\leq 4\big(|g_{2}({\mathbf{x}})|^{2}+([g_{1}({\mathbf{x}})]_{+})^{2}\big).$
If $x_{1}\leq 0$, then $g_{1}({\mathbf{x}})]_{+}=0$ and $x_{1}\in[-1/2,0]$, so $1+x_{1}\geq 1/2$ and hence
$|g_{2}({\mathbf{x}})|=|x_{1}+x_{1}^{2}|=|x_{1}|\,(1+x_{1})\geq\frac{1}{2}|x_{1}|.$
Thus,
$\text{dist}({\mathbf{x}},\mathcal{S}(I_{g},\emptyset))^{2}=|x_{1}|^{2}\leq 4|g_{2}({\mathbf{x}})|^{2}\leq 4\big(|g_{2}({\mathbf{x}})|^{2}+([g_{1}({\mathbf{x}})]_{+})^{2}\big).$
Combining Case 1 and Case 2, we conclude that for all ${\mathbf{x}}\in\mathcal{X}\cap\mathbb{B}({\mathbf{x}}^{*},\delta)$, all $I_{g}\subseteq J_{g}({\mathbf{x}}^{*})$, and all $I_{A}\subseteq J_{A}({\mathbf{x}}^{*})$ (with $J_{A}({\mathbf{x}}^{*})=\emptyset$), (14) holds with $d=1$ and $\gamma=4$. This example shows that (14) with $d=1$ may hold even when the local LICQ fails at $\mathbf{x}$$^{*}$. This completes the proof.
Appendix B Proof of Technical Results
In this section, we present the proofs of the technical results for the complexity analysis.
B.1 Technical Results for Subproblem Accuracy
The following lemma shows that, when the inexact solution ${\mathbf{x}}^{(t+1)}$ of the subproblem $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ satisfies (23), it also produces a small objective gap for this subproblem.
Lemma B.1.
Suppose ${\mathbf{x}}^{(t+1)}\in\mathcal{X}$ satisfies (23). Then it holds that
$\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\leq\frac{\epsilon_{t}^{2}}{2(p-L)}.$ (55)
Proof.
Condition (23) is equivalent to the existence of ${\mathbf{v}}^{(t+1)}\in\mathbb{R}^{n}$ and $\boldsymbol{\xi}^{(t+1)}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})$ such that $\|{\mathbf{v}}^{(t+1)}\|\leq\epsilon_{t}$ and
$\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})=-\,\boldsymbol{\xi}^{(t+1)}-{\mathbf{v}}^{(t+1)}.$ (56)
By the definition of the normal cone, this implies that for any ${\mathbf{x}}\in\mathcal{X}$,
$\langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}-{\mathbf{x}}^{(t+1)}\rangle=\langle\boldsymbol{\xi}^{(t+1)},{\mathbf{x}}^{(t+1)}-{\mathbf{x}}\rangle\geq 0.$ (57)
In addition, since $\mathcal{L}_{p}(\cdot,{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ is (pL)(p-L)-strongly convex in $\mathbf{x}$}, we have
$\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\leq$ $-\langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\rangle\textstyle-\frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}$
$\leq$ $\penalty 10000\ \langle{\mathbf{v}}^{(t+1)},\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\rangle-\frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}$
$\leq$ $\penalty 10000\ \frac{1}{2(p-L)}\|{\mathbf{v}}^{(t+1)}\|^{2}\leq\frac{\epsilon_{t}^{2}}{2(p-L)},$
where the second inequality follows from (57) with ${\mathbf{x}}=\tilde{{\mathbf{x}}}^{(t+1)}$, and the last inequality follows from Young’s inequality. The final result yields (55). ∎
B.2 Proof of Lemma 4.1
Proof of Lemma 4.1.
Let ${\mathbf{v}}^{(t+1)}$ be the vector from (56). Recall that ${\mathbf{x}}_{\text{feas}}\in\mathcal{X}$ is from Assumption 3.1F. By (57) at ${\mathbf{x}}={\mathbf{x}}_{\text{feas}}$, we have
$\langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}^{(t+1)}-{\mathbf{x}}_{\text{feas}}\rangle\leq 0.$ (58)
Using the convexity of $g_{i}$’s and $\boldsymbol{\lambda}^{(t+1)}\geq{\mathbf{0}}$, we have
$\langle\boldsymbol{\lambda}^{(t+1)},{\mathbf{g}}({\mathbf{x}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\penalty 10000\ \textstyle\sum_{i=1}^{m}\langle\lambda_{i}^{(t+1)}\nabla g_{i}({\mathbf{x}}^{(t+1)}),{\mathbf{x}}^{(t+1)}-{\mathbf{x}}_{\text{feas}}\rangle$
$=$ $\penalty 10000\ \langle\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}^{(t+1)}-{\mathbf{x}}_{\text{feas}}\rangle$
$+\langle\nabla f({\mathbf{x}}^{(t+1)})+p({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{(t+1)}\rangle$
$\leq$ $\penalty 10000\ \langle\nabla f({\mathbf{x}}^{(t+1)})+p({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})+{\mathbf{v}}^{(t+1)},{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}^{(t+1)}\rangle$
$\leq$ $\penalty 10000\ (B_{f}+pD_{\mathcal{X}}+1)\cdot D_{\mathcal{X}}=C_{\boldsymbol{\lambda}},$ (59)
where the second inequality is by (58) and the last one is from Assumption 3.1A and D, and the fact that $\|{\mathbf{v}}^{(t+1)}\|\leq\epsilon_{t}<1$. Since $\boldsymbol{\lambda}^{(0)}={\mathbf{0}}$ in Algorithm 1, (59) also holds for t=1t=-1. So, $\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq C_{\boldsymbol{\lambda}},\penalty 10000\ \forall\,t\geq 0$, or equivalently,
$\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle\leq C_{\boldsymbol{\lambda}}-\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle,\penalty 10000\ \forall\,t\geq 0.$ (60)
We then prove the result in (24) by induction. It clearly holds for $t=0$ since $\boldsymbol{\lambda}^{(0)}={\mathbf{0}}$. Suppose that the inequality in (24) holds for some $t\geq 0$. By $\|{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\|\leq B_{g}$ (see Assumption 3.1E) and $\overline{\tau}=\sup_{t\geq 0}\tau_{t}$, we have $\|\tau_{t}{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\|\leq\overline{\tau}B_{g}\leq M_{\boldsymbol{\lambda}}/2$. Since $\|\boldsymbol{\lambda}^{(t)}\|\leq M_{\boldsymbol{\lambda}}$, we have
$\|\boldsymbol{\lambda}^{(t)}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\penalty 10000\ \max_{\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|\leq M_{\boldsymbol{\lambda}}}\left\{\|\boldsymbol{\lambda}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}$
$=$ $\penalty 10000\ \max_{\theta\in[0,1],\,\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|=M_{\boldsymbol{\lambda}}}\left\{\|\theta\boldsymbol{\lambda}\|^{2}-2\theta\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}.$ (61)
Fixing any $\boldsymbol{\lambda}$ satisfying $\boldsymbol{\lambda}\geq{\mathbf{0}}$ and $\|\boldsymbol{\lambda}\|=M_{\boldsymbol{\lambda}}$, consider the maximization in (61) only over $\theta\in[0,1]$. Since $\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle/\|\boldsymbol{\lambda}\|^{2}\leq\|\tau_{t}{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\|/M_{\boldsymbol{\lambda}}\leq 1/2$, $\theta=1$ is an optimal solution for (61). As a consequence,
$\|\boldsymbol{\lambda}^{(t)}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\leq$ $\penalty 10000\ \max_{\boldsymbol{\lambda}\geq{\mathbf{0}},\,\|\boldsymbol{\lambda}\|=M_{\boldsymbol{\lambda}}}\left\{\|\boldsymbol{\lambda}\|^{2}-2\tau_{t}\langle\boldsymbol{\lambda},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\right\}$
$=$ $\penalty 10000\ \textstyle M_{\boldsymbol{\lambda}}^{2}-2\tau_{t}M_{\boldsymbol{\lambda}}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})].$ (62)
According to the updating equation of $\boldsymbol{\lambda}^{(t+1)}$ in Step 4 of Algorithm 1, we have
$\|\boldsymbol{\lambda}^{(t+1)}\|^{2}=$ $\penalty 10000\ \textstyle\sum_{i=1}^{m}([\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)})]_{+})^{2}$
$\leq$ $\penalty 10000\ \textstyle\sum_{i=1}^{m}(\lambda_{i}^{(t)}+\tau_{t}\cdot g_{i}({\mathbf{x}}^{(t)}))^{2}=\|\boldsymbol{\lambda}^{(t)}\|^{2}+2\tau_{t}\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\tau_{t}^{2}\|{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}$
$\leq$ $\penalty 10000\ \|\boldsymbol{\lambda}^{(t)}\|^{2}+2\tau_{t}\big(C_{\boldsymbol{\lambda}}-\langle\boldsymbol{\lambda}^{(t)},-{\mathbf{g}}({\mathbf{x}}_{\text{feas}})\rangle\big)+\tau_{t}^{2}\|{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}$
$\leq$ $\penalty 10000\ \textstyle M_{\boldsymbol{\lambda}}^{2}+2\overline{\tau}C_{\boldsymbol{\lambda}}-2\underline{\tau}M_{\boldsymbol{\lambda}}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]+\overline{\tau}^{2}B_{g}^{2}\leq M_{\boldsymbol{\lambda}}^{2},$
where the second inequality is from (60), and the last one is from (62), the definitions of $\overline{\tau}$, $\underline{\tau}$ and $M_{\boldsymbol{\lambda}}$ in the settings. (24) then follows by the inductive result. ∎
B.3 Proof of Lemma 4.2
Proof of Lemma 4.2.
Recall that function $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ is (pL)(p-L)-strongly convex and $K$-smooth in $\mathbf{x}$} over $\mathcal{X}$ for any $t\geq 0$. Suppose the APG method by Nesterov [37, Eq. (2.2.63)] is applied to $\min_{{\mathbf{x}}\in\mathcal{X}}\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$ for $k_{t}$ iterations with any initial solution in ${\mathbf{u}}^{(0)}\in\mathcal{X}$ and returns a solution ${\mathbf{u}}^{(k_{t})}$. According to Nesterov [37, Theorem 2.2.1–2.2.3 & Lemma 2.2.4], Lemma 4.1, and Assumption 3.1A, D, and E, we have
$\penalty 10000\ \mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\leq$ $\left(1-\sqrt{\frac{p-L}{K}}\right)^{k_{t}}\left(\mathcal{L}_{p}({\mathbf{u}}^{(0)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+\frac{p-L}{2}\|{\mathbf{u}}^{(0)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\right)$
$\leq$ $\left(1-\sqrt{\frac{p-L}{K}}\right)^{k_{t}}\left((B_{f}+M_{\boldsymbol{\lambda}}B_{g})D_{\mathcal{X}}+\frac{2p-L}{2}D_{\mathcal{X}}^{2}\right).$
Hence, it must hold that
$\mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\leq\frac{\epsilon_{t}^{2}}{8K}$ (63)
for some $k_{t}$ satisfying
$k_{t}=$ $\penalty 10000\ \sqrt{\frac{K}{p-L}}\ln\left(\frac{(B_{f}+M_{\boldsymbol{\lambda}}B_{g})D_{\mathcal{X}}+\frac{2p-L}{2}D_{\mathcal{X}}^{2}}{\epsilon_{t}^{2}/(8K)}\right)$
$=$ $\penalty 10000\ O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right).$ (64)
By the definition of ${\mathbf{x}}^{(t+1)}$, we have
$\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})+K({\mathbf{x}}^{(t+1)}-{\mathbf{u}}^{(k_{t})})+{\mathbf{v}}^{(t+1)}=0$ (65)
for some ${\mathbf{v}}^{(t+1)}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})$. Finally, according to Nesterov [36, Theorem 1],
$\text{dist}\big(-\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)$
$\leq$ $\penalty 10000\ 2K\|{\mathbf{u}}^{(k_{t})}-{\mathbf{x}}^{(t+1)}\|$
$\leq$ $\penalty 10000\ 2\sqrt{2K\left(\mathcal{L}_{p}({\mathbf{u}}^{(k_{t})},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\right)}\leq\epsilon_{t},$
where the first inequality is because of (65) and the $K$-smoothness of $\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$, the second inequality is by Nesterov [36, Theorem 1], and the last one is because of (63). This means (23) is ensured with complexity $O(\ln(\epsilon_{t}^{-1}))$ for any $t\geq 0$. ∎
B.4 Proof of Proposition 5.2
To prove Proposition 5.2, we need to first present three descent lemmas similar to the ones in Zhang and Luo [53, 54], Zhang et al. [55].
Lemma B.2.
The sequence $\{({\mathbf{x}}^{(t)},\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$\geq$ $\penalty 10000\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle$
$+\frac{p}{2}(2/\theta_{t}-1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}.$ (66)
Proof.
According to the definition of $\mathcal{L}_{p}$ in (19), we have
$\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})=-\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle.$ (67)
By the (pL)(p-L)-strong convexity of $\mathcal{L}_{p}$ in $\mathbf{x}$}, we have
$\mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\geq\frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}.$ (68)
Moreover, we have from Step 6 of Algorithm 1 that
$\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\penalty 10000\ (p/2)\cdot(\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}-\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t+1)}\|^{2})$
$=$ $\penalty 10000\ (p/2)\cdot\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)})+({\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t+1)})\rangle$
$=$ $\penalty 10000\ (p/2)\cdot(2/\theta_{t}-1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$ (69)
for $\theta_{t}\leq 1$. Combining (55), (67), (68) and (69) yields (66). ∎
Lemma B.3.
The sequence $\{(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})\geq$ $\penalty 10000\ \langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle$
$+\frac{p}{2}\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{z}}^{(t+1)}+{\mathbf{z}}^{(t)}-2{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})\rangle.$ (70)
Proof.
By the definition of $\mathcal{L}_{p}$ in (19) and the definition of $d$ in (20), we have
$d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})=$ $\penalty 10000\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})$
$\geq$ $\penalty 10000\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})$
$=$ $\penalty 10000\ \langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle.$
Next, using the same technique, we have
$\penalty 10000\ d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$
$=$ $\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\geq$ $\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\penalty 10000\ \frac{p}{2}(\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-{\mathbf{z}}^{(t+1)}\|^{2}-\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-{\mathbf{z}}^{(t)}\|^{2})$
$=$ $\penalty 10000\ \frac{p}{2}\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{z}}^{(t+1)}+{\mathbf{z}}^{(t)}-2{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})\rangle.$
Combining the above inequalities yields (70). ∎
Lemma B.4.
The sequence $\{{\mathbf{z}}^{(t)}\}_{t\geq 0}$ generated by Algorithm 1 satisfies that
$v({\mathbf{z}}^{(t+1)})-v({\mathbf{z}}^{(t)})\leq p\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{z}}^{(t)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\rangle+\frac{p(1/\sigma+1)}{2}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}.$ (71)
Proof.
According to (18) and Lemma 5.1, for any $\mathbf{z}$} and $\mathbf{z}$$^{\prime}$ in $\mathcal{X}$, it holds that
$\|\nabla v({\mathbf{z}})-\nabla v({\mathbf{z}}^{\prime})\|=$ $\penalty 10000\ \|p({\mathbf{z}}-{\mathbf{x}}({\mathbf{z}}))-p({\mathbf{z}}^{\prime}-{\mathbf{x}}({\mathbf{z}}^{\prime}))\|$
$\leq$ $\penalty 10000\ p(\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|+\|{\mathbf{x}}({\mathbf{z}})-{\mathbf{x}}({\mathbf{z}}^{\prime})\|)$
$\leq$ $\penalty 10000\ p(1/\sigma+1)\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|,$
meaning that $v({\mathbf{z}})$ is $p(1/\sigma+1)$-smooth, which implies (71). ∎
Using the three descent lemmas above, Proposition 5.2 can be proved as follows.
Proof Proposition 5.2.
Applying (66), (70) and (71) to (28) and rearranging terms, we have
$\phi^{t}-\phi^{t+1}=$ $\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$-2(d(\boldsymbol{\lambda}^{(t)},{\mathbf{z}}^{(t)})-d(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)}))+2(v({\mathbf{z}}^{(t)})-v({\mathbf{z}}^{(t+1)}))$
$\geq$ $\penalty 10000\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}+\frac{p}{2}(2/\theta_{t}-1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$-\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle$
$+p\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}-2({\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-{\mathbf{x}}({\mathbf{z}}^{(t)}))\rangle$
$-p(1/\sigma+1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}.$ (72)
For any $\alpha>0$, by Young’s inequality, it holds that
$2\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\rangle\leq\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}/\alpha+\alpha\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}.$ (73)
By Cauchy-Schwartz inequality and (26), we have
$\penalty 10000\ 2\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-\tilde{{\mathbf{x}}}^{(t+1)}\rangle$
$\leq$ $\penalty 10000\ 2\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|\cdot\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-\tilde{{\mathbf{x}}}^{(t+1)}\|\leq(2/\sigma)\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}.$ (74)
Applying (73) and (74) to (72), we have
$\phi^{t}-\phi^{t+1}\geq$ $\penalty 10000\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}-p\alpha\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$+p(1/\theta_{t}-{3/2}-1/\alpha-3/\sigma)\cdot\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$+2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle.$ (75)
According to the updating equation of $\boldsymbol{\lambda}^{(t+1)}$ in Step 4 of Algorithm 1, we have
$-\tau_{t}\langle\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle\geq-\tau_{t}\langle\boldsymbol{\lambda}^{(t+1)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\|\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)}\|^{2},$ (76)
$-\boldsymbol{\lambda}^{(t+1)}+\boldsymbol{\lambda}^{(t)}+\tau_{t}{\mathbf{g}}({\mathbf{x}}^{(t)})\in\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}).$ (77)
Therefore, we have
$\penalty 10000\ 2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle$
$\geq$ $\penalty 10000\ 2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\tau_{t}^{-1}\cdot\|\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)}\|^{2}$
$=$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\frac{1}{\tau_{t}}\|\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)}+\tau_{t}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)}))\|^{2}$
$\geq$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\frac{1}{\tau_{t}}\cdot\text{dist}\big(\tau_{t}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}$
$=$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\tau_{t}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2},$ (78)
where the two inequalities are by (76) and (77), respectively, and the last equality is from the structure of $\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})$.
Applying (78) and Assumption 3.1E to (75), we have
$\phi^{t}-\phi^{t+1}\geq$ $\left(\frac{p-L}{2}-\tau_{t}B_{g}^{2}\right)\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}-p\alpha\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$+p(1/\theta_{t}-{3/2}-1/\alpha-3/\sigma)\cdot\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$+\tau_{t}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}.$ (79)
Let $\alpha=6\theta$. Recall that $\theta_{t}=\theta\leq\frac{p-L}{18p}<\frac{1}{18}$ and $\sigma=\frac{p-L}{p}$. We have
$\frac{1}{\theta_{t}}-{3/2}-1/\alpha-3/\sigma\geq\frac{1}{2\theta_{t}}-1/\alpha-3/\sigma=\frac{1}{3\theta}-\frac{3p}{p-L}\geq\frac{1}{6\theta}.$ (80)
Recall that $\tau_{t}=\tau=\frac{p-L}{4B_{g}^{2}}$. We have $\frac{p-L}{2}-\tau_{t}B_{g}^{2}=\frac{p-L}{4}$. Applying this equation, (80), and $\alpha=6\theta$ to (79) leads to (30). ∎
Appendix C Sufficient Decrease of Potential Function $\phi_{t}$
In this section, we present a series of technical results that establish the sufficient decrease of the potential function $\phi_{t}$ after each iteration of Algorithm 1. Such sufficient decrease is the key to establish the total complexity.
C.1 Proof of Proposition 5.3
To prove Proposition 5.3, we first need the following lemma.
Lemma C.1.
For any $\delta^{\prime}>0$, there must exist $R(\delta^{\prime})>0$ such that, if
$\textup{dist}\big(-\nabla f(\tilde{{\mathbf{x}}})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}})\boldsymbol{\lambda},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}})\big)^{2}+\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda})\big)^{2}\leq R^{2}(\delta^{\prime})$ (81)
for some $\tilde{\mathbf{x}}\in\mathcal{X}$ and some $\boldsymbol{\lambda}\geq{\mathbf{0}}$ satisfying $\|\boldsymbol{\lambda}\|\leq M_{\boldsymbol{\lambda}}$, there must exist ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|\tilde{\mathbf{x}}-{\mathbf{x}}^{*}\|\leq\delta^{\prime}$ and $\textup{supp}(\boldsymbol{\lambda})\subset J_{g}({\mathbf{x}}^{*})\text{ and }J_{A}(\tilde{\mathbf{x}})\subset J_{A}({\mathbf{x}}^{*})$.
Proof.
We prove this by contradiction. Suppose that the claim does not hold. Then there exists a sequence of $\{(\tilde{{\mathbf{x}}}^{(k)},\tilde{\boldsymbol{\lambda}}^{(k)})\}_{k\geq 1}\subset\mathcal{X}\times\mathbb{R}_{+}^{m}$, $\|\tilde{\boldsymbol{\lambda}}^{(k)}\|\leq M_{\boldsymbol{\lambda}}$, such that
$\text{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{(k)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(k)})\tilde{\boldsymbol{\lambda}}^{(k)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(k)})\big)^{2}+\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(k)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\tilde{\boldsymbol{\lambda}}^{(k)})\big)^{2}\leq 1/k,$ (82)
but for any $k$, there is no ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|\tilde{\mathbf{x}}^{(k)}-{\mathbf{x}}^{*}\|\leq\delta^{\prime}$ and $\textup{supp}(\tilde{\boldsymbol{\lambda}}^{(k)})\subset J_{g}({\mathbf{x}}^{*})$ and $J_{A}(\tilde{\mathbf{x}}^{(k)})\subset J_{A}({\mathbf{x}}^{*})$.
[…]
$\text{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{*})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{*})\tilde{\boldsymbol{\lambda}}^{*},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{*})\big)^{2}+\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{*}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\tilde{\boldsymbol{\lambda}}^{*})\big)^{2}=0,$
[…]
With Lemma C.1, we are able to prove Proposition 5.3.
Proof of Proposition 5.3.
Let $\delta$ be given by Assumption 3.3 and let $R(\cdot)$ be as in Lemma C.1. By (20), we have $\tilde{\boldsymbol{\xi}}^{(t+1)}:=-\nabla_{\mathbf{x}}\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\in\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)}).$ Since $\mathcal{L}_{p}$ is $K$-smooth in $\mathbf{x}$}, it holds for any ${\mathbf{x}}\in\mathcal{X}$ that
$\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\leq-\langle\tilde{\boldsymbol{\xi}}^{(t+1)},{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\rangle+(K/2)\cdot\|{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}.$ (83)
By the (pL)(p-L)-strongly convexity of $\mathcal{L}_{p}$ in $\mathbf{x}$}, it holds that
$\frac{p-L}{2}\|{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}).$ (84)
Adding (83) and (84) with ${\mathbf{x}}={\mathbf{x}}({\mathbf{z}}^{(t)})$ yields
$\frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$-\langle\tilde{\boldsymbol{\xi}}^{(t+1)},{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\rangle+(K/2)\cdot\|{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}.$ (85)
Recall that $\|\boldsymbol{\lambda}^{(t+1)}\|\leq M_{\boldsymbol{\lambda}}$. According to Lemma C.1 and (31), there must exist ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|\tilde{\mathbf{x}}^{(t+1)}-{\mathbf{x}}^{*}\|\leq\delta$, $I_{g}^{(t+1)}=\text{supp}(\boldsymbol{\lambda}^{(t+1)})\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}^{(t+1)}=J_{A}(\tilde{\mathbf{x}}^{(t+1)})\subset J_{A}({\mathbf{x}}^{*})$. Let $\bar{{\mathbf{x}}}^{(t+1)}$ be the projection of $\tilde{\mathbf{x}}^{(t+1)}$ onto $\mathcal{S}(I_{g}^{(t+1)},I_{A}^{(t+1)})$, so that $\text{dist}(\tilde{\mathbf{x}}^{(t+1)},\mathcal{S}(I_{g}^{(t+1)},I_{A}^{(t+1)}))=\|\tilde{\mathbf{x}}^{(t+1)}-\bar{{\mathbf{x}}}^{(t+1)}\|$. As a result, it holds that
$\penalty 10000\ g_{i}(\bar{{\mathbf{x}}}^{(t+1)})=0,\;i\in I_{g}^{(t+1)},\;g_{i}(\bar{{\mathbf{x}}}^{(t+1)})\leq 0,\;i\in[m]\backslash I_{g}^{(t+1)},$ (86)
[…]
Applying Assumption 3.3 and using $I_{A}^{(t+1)}=J_{A}(\tilde{\mathbf{x}}^{(t+1)})$, we obtain
$\|\bar{{\mathbf{x}}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2d}\leq$ $\penalty 10000\ \gamma^{2}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}$
$\iff\|\bar{{\mathbf{x}}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\penalty 10000\ \gamma^{2/d}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2/d}.$ (88)
Let ${\mathbf{A}}(I_{A}^{(t+1)},:)$ be the matrix formed by the rows of $\mathbf{A}$} with indices in $I_{A}^{(t+1)}$. Since $\tilde{\boldsymbol{\xi}}^{(t+1)}\in\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})$, there exists some ${\mathbf{w}}^{(t+1)}\in\mathbb{R}_{+}^{|I_{A}^{(t+1)}|}$ such that $\tilde{\boldsymbol{\xi}}^{(t+1)}=[{\mathbf{A}}(I_{A}^{(t+1)},:)]^{\top}{\mathbf{w}}^{(t+1)},$ which, according to (87), implies
$\langle\tilde{\boldsymbol{\xi}}^{(t+1)},\bar{{\mathbf{x}}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\rangle=0.$ (89)
Setting ${\mathbf{x}}=\bar{{\mathbf{x}}}^{(t+1)}$ in (85) and applying (88) and (89), we have
$\frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\bar{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$+(\gamma^{1/d}K/2)\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2/d}.$ (90)
By (17) and (86), we must have $\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)}))\leq 0$ and $\lambda_{i}^{(t+1)}g_{i}(\bar{{\mathbf{x}}}^{(t+1)})=0$ for $i\in[m]$. This implies
$\penalty 10000\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\bar{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\leq$ $\left(f({\mathbf{x}}({\mathbf{z}}^{(t)}))+\frac{p}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}\|^{2}\right)-\left(f(\bar{{\mathbf{x}}}^{(t+1)})+\frac{p}{2}\|\bar{{\mathbf{x}}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\right)\leq 0,$ (91)
where the last inequality is because $\bar{{\mathbf{x}}}^{(t+1)}$ is a feasible solution to the minimization problem in (17) while ${\mathbf{x}}({\mathbf{z}}^{(t)})$ is its optimal solution. Combining (91) with (90) yields the claim. ∎
C.2 Proof of Lemma 5.4
Proof of Lemma 5.4.
By the (pL)(p-L)-strong convexity of $\mathcal{L}_{p}$ in $\mathbf{x}$}, we have
$\penalty 10000\ \frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),$
$\penalty 10000\ \frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)}))-\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)})),$
where $\boldsymbol{\lambda}({\mathbf{z}}^{(t)})$ is a vector of Lagrangian multipliers corresponding to ${\mathbf{x}}({\mathbf{z}}^{(t)})$ defined in (17). Notice that by the strong duality (see, e.g. Rockafellar [41, Sec. 28–30]), we have $\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\in\mathop{\mathrm{arg\,max}{}}_{\boldsymbol{\lambda}\geq{\mathbf{0}}}\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda})$. Hence, adding the above two inequalities gives
$(p-L)\cdot\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq$ $\penalty 10000\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)}))-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\penalty 10000\ \langle\boldsymbol{\lambda}({\mathbf{z}}^{(t)})-\boldsymbol{\lambda}^{(t+1)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle$
$\leq$ $\penalty 10000\ \textstyle\sum_{i\in\mathrm{supp}(\boldsymbol{\lambda}^{(t+1)})}|\lambda_{i}({\mathbf{z}}^{(t)})-\lambda_{i}^{(t+1)}|\cdot|g_{i}(\tilde{{\mathbf{x}}}^{(t+1)})|$
[…]
$\leq$ $\penalty 10000\ \|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})-\boldsymbol{\lambda}^{(t+1)}\|\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big),$ (92)
where the last inequality follows from the Cauchy–Schwarz inequality.
In addition, by the KKT conditions at $({\mathbf{x}}({\mathbf{z}}^{(t)}),\boldsymbol{\lambda}({\mathbf{z}}^{(t)}))$ in (17), we have
$\textstyle\boldsymbol{\xi}:=-\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))-\sum_{i=1}^{m}\lambda_{i}({\mathbf{z}}^{(t)})\nabla g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)}))-p({\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)})\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}({\mathbf{z}}^{(t)})),$ (93)
$\lambda_{i}({\mathbf{z}}^{(t)})g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)}))=0,\;i\in[m].$ (94)
Hence, by the convexity of $g_{i}$’s and $\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\geq{\mathbf{0}}$, it follows that
$\textstyle\sum_{i=1}^{m}\lambda_{i}({\mathbf{z}}^{(t)})\cdot g_{i}({\mathbf{x}}_{\text{feas}})\quad\geq\;\;\;\;$ $\penalty 10000\ \textstyle\sum_{i=1}^{m}\lambda_{i}({\mathbf{z}}^{(t)})\left(g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)}))+\big\langle\nabla g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)})),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\big\rangle\right)$
[…]
$\qquad\geq\;\;\;\;$ $\penalty 10000\ \big\langle\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))+p({\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}),{\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{x}}_{\text{feas}}\big\rangle,$
where ${\mathbf{x}}_{\text{feas}}$ is given in Assumption 3.1F, and the last inequality holds because $\boldsymbol{\xi}\in\mathcal{N}_{\mathcal{X}}({\mathbf{x}}({\mathbf{z}}^{(t)}))$. Now applying Assumption 3.1A, D, and F, we obtain
$\penalty 10000\ \textstyle\|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\|_{1}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]$
$\leq$ $\penalty 10000\ \langle\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))+p({\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\rangle$
$\leq$ $\penalty 10000\ \big(\|\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))\|+p\cdot\|{\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}\|\big)\cdot\|{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|$
$\leq$ $\penalty 10000\ (B_{f}+pD_{\mathcal{X}})\cdot D_{\mathcal{X}},$
and thus $\|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\|\leq\|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\|_{1}\leq\frac{(B_{f}+pD_{\mathcal{X}})\cdot D_{\mathcal{X}}}{\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]}\leq M_{\boldsymbol{\lambda}},$ which together with (92) and (24) gives the desired result. ∎
C.3 Proof of Proposition 5.5
Proof of Proposition 5.5.
By the choices of $\tau_{t}$ and $\theta_{t}$, inequality (30) holds by Proposition 5.2. We then prove (36) by considering two cases depending on whether (31) holds.
Case I: (31) is satisfied. In this case, inequality (32) holds by Proposition 5.3. Let
$s:=\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big),\quad C_{d}:=\frac{\gamma^{1/d}K}{p-L}.$
Then (32) implies
$\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq C_{d}s^{2/d},$
and hence
$\penalty 10000\ \tau\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$\geq$ $\penalty 10000\ \tau s^{2}-6p\theta C_{d}s^{2/d}.$
When $d\in(1,2]$, let
$r:=\frac{d}{d-1},\quad r^{\prime}:=d,\quad\frac{1}{r}+\frac{1}{r^{\prime}}=1.$
We apply Young’s inequality in the form
$AB\leq A^{r}/r+\frac{B^{r^{\prime}}}{r^{\prime}},\quad A,B\geq 0.$
Choose
$A:=\left(\frac{2}{\tau d}\right)^{1/d}6p\theta C_{d},\quad B:=\left(\frac{\tau d}{2}\right)^{1/d}s^{2/d}.$
Then
$6p\theta C_{d}s^{2/d}=AB\leq$ $\penalty 10000\ \frac{1}{r}\left(\left(\frac{2}{\tau d}\right)^{1/d}6p\theta C_{d}\right)^{r}+\frac{1}{r^{\prime}}\left(\left(\frac{\tau d}{2}\right)^{1/d}s^{2/d}\right)^{r^{\prime}}$
$=$ $\penalty 10000\ \left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta C_{d}\big)^{\frac{d}{d-1}}+\frac{\tau}{2}s^{2},$
where the equality uses $\frac{1}{r}=1-\frac{1}{d}$ and $\frac{1}{r^{\prime}}=\frac{1}{d}$. Rearranging yields
$\tau s^{2}-6p\theta C_{d}s^{2/d}\geq$ $\penalty 10000\ \frac{\tau}{2}s^{2}-\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta C_{d}\big)^{\frac{d}{d-1}}$
$=$ $\penalty 10000\ \frac{\tau}{2}s^{2}-\Xi_{d}(\theta),$
where $\Xi_{d}(\theta)$ is defined in (37).
When $d=1$, inequality (32) reduces to the linear bound
$\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq C_{1}s^{2},\penalty 10000\ \text{ with }\penalty 10000\ C_{1}:=\frac{\gamma K}{p-L}.$
Hence
$\tau s^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\geq(\tau-6p\theta C_{1})s^{2}.$
By the definition of $\hat{\theta}_{d}$ with $d=1$, we have $\theta\leq\hat{\theta}_{d}=\frac{\tau}{12pC_{1}}$, which yields
[…]
Hence,
$\penalty 10000\ \tau\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$\geq$ $\penalty 10000\ \frac{\tau}{2}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2},$
which corresponds to $\Xi_{1}(\theta)=0$ given in (37).
In both subcases,
$\penalty 10000\ \tau\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
$\geq$ $\penalty 10000\ \frac{\tau}{2}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\Xi_{d}(\theta).$ (95)
Then applying (95) to (30) implies (36).
Case II: Condition (31) does not hold. In this case, it holds that
$\text{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big)^{2}$ (96)
$+\;\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}$ $>R^{2}(\delta).$
By the triangle inequality and $\nabla_{\mathbf{x}}\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\in\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})$, it holds that
$\penalty 10000\ \text{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big)^{2}$
$\leq$ $\penalty 10000\ p^{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\leq 2p^{2}\big(\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}+\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\big)$
$=$ $\penalty 10000\ 2p^{2}\big(\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}^{(t+1)}\|^{2}+(1/\theta^{2})\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\big)$
$\leq$ $\penalty 10000\ 2p^{2}\left(\frac{\epsilon_{t}^{2}}{(p-L)^{2}}+\frac{1}{\theta^{2}}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\right),$ (97)
where the last inequality follows from (55) and (84) with ${\mathbf{x}}={\mathbf{x}}^{(t+1)}$. Plugging (97) into (96), we obtain
$\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}+2p^{2}\left(\frac{\epsilon_{t}^{2}}{(p-L)^{2}}+\frac{1}{\theta^{2}}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}\right)>R^{2}(\delta).$ (98)
We then have
$\penalty 10000\ \frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
[…]
$\geq\,$ $\penalty 10000\ \frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\frac{3\tau}{4}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\frac{1}{\tau}\left(\frac{12p\theta M_{\boldsymbol{\lambda}}}{p-L}\right)^{2}$
$\geq\,$ $\penalty 10000\ \frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\frac{\tau}{2}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\frac{1}{\tau}\left(\frac{12p\theta M_{\boldsymbol{\lambda}}}{p-L}\right)^{2}+\frac{R^{2}(\delta)\theta}{24p}-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}}$
$\geq\,$ $\penalty 10000\ \frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\frac{\tau}{2}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\frac{p\theta\epsilon_{t}^{2}}{12(p-L)^{2}},$ (99)
where the second inequality follows from the Young’s inequality, the third one holds by (98) and $\theta\leq 6p\tau$, and the last one results from $\theta\leq\frac{R^{2}(\delta)(p-L)^{2}\tau}{2\cdot 12^{3}p^{3}M_{\boldsymbol{\lambda}}^{2}}$. Applying (99) to (30) gives (36). ∎
Appendix D Proof of Theorem 5.6
Proof of Theorem 5.6.
Consider any $t\in\{0,1,\dots,T-1\}$. Firstly, by (23) and the triangle inequality, we have
$\penalty 10000\ \text{dist}\big(-\nabla f({\mathbf{x}}^{(t+1)})-\textstyle\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\cdot\nabla g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)$
$\leq$ $\penalty 10000\ \text{dist}\big(-\nabla_{\mathbf{x}}\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)+p\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|$
$\leq$ $\penalty 10000\ \epsilon_{t}+p\|{\mathbf{x}}^{(t+1)}-{\mathbf{z}}^{(t)}\|=\epsilon_{t}+(p/\theta)\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|.$ (100)
Secondly, recalling that $\tilde{{\mathbf{x}}}^{(t+1)}={\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t)})$, we have
$\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|\leq$ $\penalty 10000\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}-[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|$
$\leq$ $\penalty 10000\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+B_{g}\|{\mathbf{x}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\|$
$\leq$ $\penalty 10000\ \|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|+\frac{B_{g}\epsilon_{t}}{p-L}\leq\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)+\frac{B_{g}\epsilon_{t}}{p-L},$ (101)
where the second inequality is by Assumption 3.1E, the third one is from (55) and (84) with ${\mathbf{x}}={\mathbf{x}}^{(t+1)}$, and the last one holds by $\|[{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}\|=\text{dist}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathbb{R}_{-}^{m})$ and $\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\subset\mathbb{R}_{-}^{m}$.
Lastly, recalling that $I_{g}^{(t+1)}=\text{supp}(\boldsymbol{\lambda}^{(t+1)})$, we have
$\textstyle\sum_{i=1}^{m}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|=$ $\penalty 10000\ \textstyle\sum_{i\in I_{g}^{(t+1)}}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|$
$\leq$ $\penalty 10000\ \textstyle\sqrt{\sum_{i\in I_{g}^{(t+1)}}(\lambda_{i}^{(t+1)})^{2}}\sqrt{\sum_{i\in I_{g}^{(t+1)}}g^{2}_{i}({\mathbf{x}}^{(t+1)})}$
$\leq$ $\penalty 10000\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big({\mathbf{g}}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)$
$\leq$ $\penalty 10000\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)+M_{\boldsymbol{\lambda}}\cdot\|{\mathbf{g}}({\mathbf{x}}^{(t+1)})-{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\|$
$\leq$ $\penalty 10000\ M_{\boldsymbol{\lambda}}\cdot\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)+\frac{M_{\boldsymbol{\lambda}}B_{g}\epsilon_{t}}{p-L},$ (102)
where the first inequality is by the Cauchy–Schwarz inequality, the second one is by Lemma 4.1 and the fact that $g^{2}_{i}({\mathbf{x}}^{(t+1)})=\text{dist}(g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}}(\lambda_{i}^{(t+1)}))^{2}$ for $i\in I_{g}^{(t+1)}$, the third one is by the triangle inequality, and the last one is by (55), (84) with ${\mathbf{x}}={\mathbf{x}}^{(t+1)}$, and Assumption 3.1E.
By summing up (36) for $t=0,1,\dots,T-1$, we have
$\penalty 10000\ \sum_{t=0}^{T-1}\left(\frac{p-L}{4}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}+\frac{p}{12\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\frac{\tau}{2}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}\right)$
$\leq$ $\penalty 10000\ \phi^{0}-\phi^{T}+\frac{p\theta}{12(p-L)^{2}}\sum_{t=0}^{T-1}\epsilon_{t}^{2}+T\cdot\Xi_{d}(\theta)\leq\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}+T\cdot\Xi_{d}(\theta),$ (103)
where the second inequality is by (29) and the fact that
$\textstyle\sum_{t=0}^{T-1}\epsilon_{t}^{2}\leq\sum_{t=0}^{\infty}\epsilon_{t}^{2}=c\pi^{2}/6.$ (104)
According to Definition 3.2, denote the squared KKT residual measured at ${\mathbf{x}}^{(t+1)}$ by
$\mathcal{R}_{t+1}:=\left(\begin{array}[]{l}\text{dist}\big(-\nabla f({\mathbf{x}}^{(t+1)})-\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\nabla g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)^{2}\\ +\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|^{2}+\big(\sum_{i=1}^{m}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|\big)^{2}\end{array}\right).$
Squaring and summing up both sides of (100), the right inequality in (101) and (102) for $t=0,1,\dots,T-1$ leads to
$\sum_{t=0}^{T-1}\mathcal{R}_{t+1}=$ $\penalty 10000\ \sum_{t=0}^{T-1}\left(\begin{array}[]{l}\text{dist}\big(-\nabla f({\mathbf{x}}^{(t+1)})-\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\nabla g_{i}({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\big)^{2}\\ +\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|^{2}+\big(\sum_{i=1}^{m}|\lambda_{i}^{(t+1)}g_{i}({\mathbf{x}}^{(t+1)})|\big)^{2}\end{array}\right)$ (107)
$\leq$ $\penalty 10000\ \frac{2p^{2}}{\theta^{2}}\sum_{t=0}^{T-1}\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}+\left(2+\frac{2(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\sum_{t=0}^{T-1}\epsilon_{t}^{2}$
$+2(M_{\boldsymbol{\lambda}}^{2}+1)\sum_{t=0}^{T-1}\text{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}$
$\leq$ $\left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\}\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}+T\cdot\Xi_{d}(\theta)\right)$
$=$ $\penalty 10000\ \mathcal{C}_{0}(\theta)+\mathcal{C}_{1}(\theta)T\cdot\Xi_{d}(\theta),$ (108)
where the second inequality is because of (103) and (104), and the equality is by the definitions of $\mathcal{C}_{0}(\theta)$ and $\mathcal{C}_{1}(\theta)$ in (38) as follows:
$\begin{aligned} \mathcal{C}_{0}(\theta):=&\penalty 10000\ \left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\}\left(\phi^{0}-\underline{f}+\frac{cp\theta\pi^{2}}{72(p-L)^{2}}\right),\\ \mathcal{C}_{1}(\theta):=&\penalty 10000\ \max\left\{\frac{24p}{\theta},\frac{4(M_{\boldsymbol{\lambda}}^{2}+1)}{\tau}\right\}.\end{aligned}$
From now on, we specify the choice of $\theta$ and the complexity of $T$ separately for $d\in(1,2]$ and $d=1$.
Firstly, we consider the case $d\in(1,2]$. Fix $\epsilon>0$. Recall from Proposition 5.5 and the definition of $\tilde{\theta}_{d}$ in (34) that, for any constant $\theta$ satisfying $0<\theta\leq\tilde{\theta}_{d}$, the sequence generated by Algorithm 1 satisfies
$\sum_{t=0}^{T-1}\mathcal{R}_{t+1}\leq\mathcal{C}_{0}(\theta)+\mathcal{C}_{1}(\theta)T\cdot\Xi_{d}(\theta)$
according to (108). We then specify
$\theta=\theta(\epsilon):=$ $\penalty 10000\ \min\left\{\tilde{\theta}_{d},\frac{6p\tau}{M_{\boldsymbol{\lambda}}^{2}+1},\left(\frac{\epsilon^{2}}{2A_{d}}\right)^{d-1}\right\},$
$\text{ where }A_{d}:=$ $\penalty 10000\ 24p\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}(6pC_{d})^{\frac{d}{d-1}}.$
Clearly, $\theta(\epsilon)\leq\tilde{\theta}_{d}$. Moreover, since $\theta(\epsilon)\leq\frac{6p\tau}{M_{\boldsymbol{\lambda}}^{2}+1}$, we have
[…]
Recalling the definition of $\Xi_{d}(\theta)$ in (37), for $d\in(1,2]$ we have
$\Xi_{d}(\theta)=\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta C_{d}\big)^{\frac{d}{d-1}},$
which implies
$\mathcal{C}_{1}(\theta(\epsilon))\cdot\Xi_{d}(\theta(\epsilon))=$ $\penalty 10000\ \frac{24p}{\theta(\epsilon)}\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}\big(6p\theta(\epsilon)\,C_{d}\big)^{\frac{d}{d-1}}$
$=$ $\penalty 10000\ 24p\left(1-\frac{1}{d}\right)\left(\frac{2}{\tau d}\right)^{\frac{1}{d-1}}(6pC_{d})^{\frac{d}{d-1}}\cdot\theta(\epsilon)^{\frac{1}{d-1}}=A_{d}\cdot\theta(\epsilon)^{\frac{1}{d-1}}.$
By the definition of $\theta(\epsilon)$, we also have
$\theta(\epsilon)\leq\left(\frac{\epsilon^{2}}{2A_{d}}\right)^{d-1},\penalty 10000\ \text{ hence }\penalty 10000\ \theta(\epsilon)^{\frac{1}{d-1}}\leq\frac{\epsilon^{2}}{2A_{d}}.$ (110)
Substituting (110) into the previous result gives
$\mathcal{C}_{1}(\theta(\epsilon))\cdot\Xi_{d}(\theta(\epsilon))=A_{d}\cdot\theta(\epsilon)^{\frac{1}{d-1}}\leq A_{d}\cdot\frac{\epsilon^{2}}{2A_{d}}=\frac{\epsilon^{2}}{2}.$ (111)
Next, denote
$B_{0}:=\left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\frac{24cp^{2}\pi^{2}}{72(p-L)^{2}},\quad B_{1}:=B_{0}+24p\big(\phi^{0}-\underline{f}\big),$ (112)
which are independent of $\epsilon$. Recalling from (38), we obtain
$\mathcal{C}_{0}(\theta(\epsilon))=$ $\penalty 10000\ \left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\mathcal{C}_{1}(\theta(\epsilon))\left(\phi^{0}-\underline{f}+\frac{cp\theta(\epsilon)\pi^{2}}{72(p-L)^{2}}\right)$
$=$ $\penalty 10000\ \left(1+\frac{(M_{\boldsymbol{\lambda}}^{2}+1)B_{g}^{2}}{(p-L)^{2}}\right)\frac{c\pi^{2}}{3}+\frac{24p}{\theta(\epsilon)}\left(\phi^{0}-\underline{f}+\frac{cp\theta(\epsilon)\pi^{2}}{72(p-L)^{2}}\right)$
$=$ $\penalty 10000\ B_{0}+\frac{24p(\phi^{0}-\underline{f})}{\theta(\epsilon)}\leq\frac{B_{1}}{\theta(\epsilon)},$ (113)
where the second equality is by (109), the third equality is by (112), and the last inequality is by (112) and the fact that $0<\theta(\epsilon)\leq 1$. We next bound $1/\theta(\epsilon)$ using the concrete choice of $\theta(\epsilon)$. Denote
$\overline{\theta}_{d}:=\min\left\{\tilde{\theta}_{d},\ \frac{6p\tau}{M_{\boldsymbol{\lambda}}^{2}+1}\right\}>0.$
If $\big(\frac{\epsilon^{2}}{2A_{d}}\big)^{d-1}\leq\overline{\theta}_{d}$, then $\theta(\epsilon)=\big(\frac{\epsilon^{2}}{2A_{d}}\big)^{d-1}$ and hence
$\frac{1}{\theta(\epsilon)}=\left(\frac{2A_{d}}{\epsilon^{2}}\right)^{d-1}.$
Otherwise, $\theta(\epsilon)=\overline{\theta}_{d}$ and thus
$\frac{1}{\theta(\epsilon)}=\frac{1}{\overline{\theta}_{d}}\leq\frac{1}{\overline{\theta}_{d}}\left(\frac{1}{\epsilon^{2}}\right)^{d-1},$
where the inequality holds since $\epsilon\in(0,1)$ and $d>1$ imply $(1/\epsilon^{2})^{d-1}\geq 1$. Thus, we conclude that
$\frac{1}{\theta(\epsilon)}\leq\frac{\max\{1/\overline{\theta}_{d},(2A_{d})^{d-1}\}}{\epsilon^{2(d-1)}},\quad\mathcal{C}_{0}(\theta(\epsilon))\leq\frac{B_{1}\max\{1/\overline{\theta}_{d},(2A_{d})^{d-1}\}}{\epsilon^{2(d-1)}}.$ (114)
Hence, we obtain
$T:=\left\lceil\frac{2\mathcal{C}_{0}(\theta(\epsilon))}{\epsilon^{2}}\right\rceil\leq\left\lceil\frac{2B_{1}\max\{1/\overline{\theta}_{d},(2A_{d})^{d-1}\}}{\epsilon^{2(d-1)}}\Big/\epsilon^{2}\right\rceil=O(\epsilon^{-2d}),$
where the inequality is by (113) and (114). Substituting $\theta=\theta(\epsilon)$ into (108), dividing both sides by $T$, and using the definition of $T$ together with (111) yield
$\frac{1}{T}\sum_{t=0}^{T-1}\mathcal{R}_{t+1}\leq\frac{\mathcal{C}_{0}(\theta(\epsilon))}{T}+\mathcal{C}_{1}(\theta(\epsilon))\,\Xi_{d}(\theta(\epsilon))\leq\frac{\epsilon^{2}}{2}+\frac{\epsilon^{2}}{2}=\epsilon^{2}.$
Therefore, there exists $s\in\{0,1,\dots,T-1\}$ such that $\mathcal{R}_{s+1}\leq\epsilon^{2}$, namely,
$\penalty 10000\ \text{dist}\big(-\nabla f({\mathbf{x}}^{(s+1)})-\textstyle\sum_{i=1}^{m}\lambda_{i}^{(s+1)}\cdot\nabla g_{i}({\mathbf{x}}^{(s+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(s+1)})\big)^{2}$ (115)
$+\|[{\mathbf{g}}({\mathbf{x}}^{(s+1)})]_{+}\|^{2}+\big(\textstyle\sum_{i=1}^{m}|\lambda_{i}^{(s+1)}g_{i}({\mathbf{x}}^{(s+1)})|\big)^{2}\penalty 10000\ \leq\penalty 10000\ \epsilon^{2}.$
According to Definition 3.2, ${\mathbf{x}}^{(s+1)}$ is an $\epsilon$-KKT point of (1).
Secondly, we consider the case $d=1$. Fix $\epsilon>0$. By setting $d=1$ in Proposition 5.5, we have $\Xi_{1}(\theta)=0$ by (37). Moreover, $C_{1}=\frac{\gamma^{1}K}{p-L}=\frac{\gamma K}{p-L}$, and the quantity $\hat{\theta}_{d}$ defined in (35) simplifies to
$\hat{\theta}_{1}=\frac{\tau}{12pC_{1}}\left(\frac{\tau}{12p}\right)^{1-1}=\frac{\tau}{12pC_{1}}=\frac{(p-L)\tau}{12p\gamma K}.$
Consequently, the admissible constant stepsize upper bound $\tilde{\theta}_{d}$ in (34) specializes to
$\tilde{\theta}_{1}=\min\left\{\frac{p-L}{18p},6p\tau,\frac{R^{2}(\delta)(p-L)^{2}\tau}{2\cdot 12^{3}p^{3}M_{\boldsymbol{\lambda}}^{2}},\frac{(p-L)\tau}{12p\gamma K}\right\},\penalty 10000\ \text{ with }\penalty 10000\ \tau=\frac{p-L}{4B_{g}^{2}}.$
Fix any constant $\theta$ satisfying $0<\theta\leq\tilde{\theta}_{1}$, and set $\theta_{t}=\theta$ for all $t\geq 0$. Then the bound (108) reduces to
$\sum_{t=0}^{T-1}\mathcal{R}_{t+1}\leq\mathcal{C}_{0}(\theta),$
where $\mathcal{C}_{0}(\theta)$ is independent of $\epsilon$. Analogous to the case $d\in(1,2]$, choosing
$T:=\left\lceil\frac{2\mathcal{C}_{0}(\theta)}{\epsilon^{2}}\right\rceil=O(\epsilon^{-2})$
ensures that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathcal{R}_{t+1}\leq\frac{\mathcal{C}_{0}(\theta)}{T}\leq\frac{\epsilon^{2}}{2}\leq\epsilon^{2}.$
Therefore, there exists $s\in\{0,1,\dots,T-1\}$ such that $\mathcal{R}_{s+1}\leq\epsilon^{2}$, and hence ${\mathbf{x}}^{(s+1)}$ is an $\epsilon$-KKT point of (1). In this case, the total number of outer iterations satisfies $T=O(\epsilon^{-2})=O(\epsilon^{-2d})$ with $d=1$.
By (64), the complexity of the $t$-th iteration of Algorithm 1 is $O\left(\sqrt{\frac{K}{p-L}}\ln(\epsilon_{t}^{-1})\right)=O(\ln(\epsilon^{-1}))$ for $t=0,1,\dots,T-1$. Therefore, the total oracle complexity of Algorithm 1 for finding an $\epsilon$-KKT point is $O(\epsilon^{-2d}\ln(\epsilon^{-1}))$. ∎
Appendix E Additional Details of Numerical Experiments
In this section, we provide additional details of the numerical experiments presented in Section 6.
E.1 Details of Datasets
We solve problem (41) on three datasets: a9a [18], bank [35] and COMPAS [1]. Details about these datasets are given in Table 1. Each dataset is splited into two subsets with a ratio of 2 : 1. The larger subset serves as $\mathcal{D}$ in the constraint, while the smaller subset is further partitioned into $\mathcal{D}_{p}$ and $\mathcal{D}_{u}$ based on the binary group variable specified in Table 1.
Table 1: Dataset information. The binary group variable represents males or females in the a9a dataset, users with age within $[25,60]$ and outside $[25,60]$ in the bank dataset, and Caucasian versus non-Caucasian individuals in the COMPAS dataset.
Datasets $n$ $d$ Label Groups
a9a 48,842 123 Income Gender
Bank 41,188 54 Subscription Age
COMPAS 6,172 16 Recidivism Race
E.2 Local LICQ and EBC for the DP classification instance (41)
Proposition E.1.
Consider problem (41) with $\mathcal{X}=\{{\mathbf{x}}\in\mathbb{R}^{d}:\|{\mathbf{x}}\|_{1}\leq r\}$. Suppose
$\left\{{\mathbf{x}}\in\mathbb{R}^{d}:\mathcal{L}({\mathbf{x}})\leq\mathcal{L}^{*}+\kappa\right\}\subset\left\{{\mathbf{x}}\in\mathbb{R}^{d}:\|{\mathbf{x}}\|_{1}\leq r/2\right\}.$
Then (41) satisfies the local LICQ condition (15) at every KKT point. Consequently, Assumption 3.3 holds with exponent $d=1$.
Proof.
Let ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ be any KKT point of (41). By the hypothesis of this proposition, the set constraint $\mathcal{X}$ is inactive in a neighborhood of $\mathbf{x}$$^{*}$ and $J_{A}({\mathbf{x}}^{*})=\emptyset$ in (15).
If $\mathcal{L}({\mathbf{x}}^{*})<\mathcal{L}^{*}+\kappa$, then no inequality constraint is active at $\mathbf{x}$$^{*}$, i.e., $J_{g}({\mathbf{x}}^{*})=\emptyset$ in (15), and the LICQ condition holds trivially.
Suppose next that $\mathcal{L}({\mathbf{x}}^{*})=\mathcal{L}^{*}+\kappa$. Since the set constraint $\mathcal{X}$ is inactive, the only active constraint is $\mathcal{L}({\mathbf{x}})\leq\mathcal{L}^{*}+\kappa$. In this case, LICQ (15) reduces to the requirement that there exists $\zeta>0$ such that $\|\nabla\mathcal{L}({\mathbf{x}}^{*})\|\geq\zeta$. Because $\mathcal{L}$ is continuous and $\mathcal{X}$ is compact, the restricted level set
$\mathcal{C}_{\kappa}:=\{{\mathbf{x}}\in\mathcal{X}:\mathcal{L}({\mathbf{x}})=\mathcal{L}^{*}+\kappa\}$
is compact whenever it is nonempty. Therefore, the quantity
$\xi:=\min\big\{\|\nabla\mathcal{L}({\mathbf{x}})\|:{\mathbf{x}}\in\mathcal{X},\ \mathcal{L}({\mathbf{x}})=\mathcal{L}^{*}+\kappa\big\}$
is attained.
Moreover, we assert that $\xi>0$, since otherwise there would exist $\bar{{\mathbf{x}}}\in\mathcal{C}_{\kappa}$ satisfying
$\nabla\mathcal{L}(\bar{{\mathbf{x}}})=\mathbf{0}\penalty 10000\ \text{ and }\penalty 10000\ \mathcal{L}(\bar{{\mathbf{x}}})=\mathcal{L}^{*}+\kappa,$
which contradicts the definition of $\mathcal{L}^{*}$ as the minimal value of $\mathcal{L}$ over $\mathcal{X}$.
Therefore, the local LICQ condition (15) holds at every KKT point. The claim follows from Proposition 3.4. ∎
E.3 Implementation Details of the Methods in Comparison
Recall that, at its $t$-th iteration, the iMELa method obtains
${\mathbf{x}}^{(t+1)}\approx\mathop{\mathrm{arg\,min}{}}_{{\mathbf{u}}\in\mathcal{X}}\left\{f({\mathbf{u}})+{\textstyle\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\cdot g_{i}({\mathbf{u}})}+\frac{p}{2}\|{\mathbf{u}}-{\mathbf{z}}^{(t)}\|^{2}\right\}.$ (116)
Similarly, at its $t$-th iteration, the iPPP method approximately solves the subproblem
${\mathbf{x}}^{(t+1)}\approx\mathop{\mathrm{arg\,min}{}}_{{\mathbf{u}}\in\mathcal{X}}\left\{f({\mathbf{u}})+\frac{\rho_{t}}{2}{\textstyle\sum_{i=1}^{m}\left(\big[g_{i}({\mathbf{u}})]_{+}\right)^{2}}+\frac{p_{t}}{2}\|{\mathbf{u}}-{\mathbf{x}}^{(t)}\|^{2}\right\},$ (117)
where $\rho_{t}>0$ is the penalty parameter and $p_{t}>0$ is the proximal parameter. In addition, DPALM follows a proximal augmented Lagrangian (AL) framework with a possibly damped dual step-size. At its $t$-th iteration, DPALM approximately solves the subproblem
${\mathbf{x}}^{(t+1)}\approx\mathop{\mathrm{arg\,min}{}}_{{\mathbf{u}}\in\mathcal{X}}\left\{f({\mathbf{u}})+\frac{p}{2}\|{\mathbf{u}}-{\mathbf{x}}^{(t)}\|^{2}+\frac{1}{2\beta_{t}}{\textstyle\sum_{i=1}^{m}}\Big(\big[\lambda_{i}^{(t)}+\beta_{t}g_{i}({\mathbf{u}})\big]_{+}^{2}-\big(\lambda_{i}^{(t)}\big)^{2}\Big)\right\},$ (118)
where $p>0$ is the proximal parameter and $\beta_{t}>0$ is the AL penalty parameter. The (possibly damped) dual step-size is chosen as
$\alpha_{t}:=\min\left\{\beta_{t},\frac{v_{t}}{\|[{\mathbf{g}}({\mathbf{x}}^{(t+1)})]_{+}\|}\right\},$
and the Lagrangian multipliers are updated by
$\lambda_{i}^{(t+1)}=\lambda_{i}^{(t)}+\alpha_{t}\max\left\{-\frac{\lambda_{i}^{(t)}}{\beta_{t}},g_{i}({\mathbf{x}}^{(t+1)})\right\},\;\forall\,i\in[m].$
For a fair comparison, we apply the accelerated projected gradient (APG) method by Nesterov [37, Eq. (2.2.63)] to approximately solve (116), (117), and (118). Let $F({\mathbf{u}})$ be the objective function in (116), (117), or (118) and let ${\mathbf{u}}^{(k)}$ be the main iterate generated by the APG method. For the task $\min_{{\mathbf{u}}\in\mathcal{X}}F({\mathbf{u}})$ where $F$ is $\mu_{F}$-strongly convex and $L_{F}$-smooth in $\mathbf{u}$}, we denote $q_{F}:=\mu_{F}/L_{F}$ and then present the APG method [37, Eq. (2.2.63)] in Algorithm 2. The APG method is terminated when
$\|\mathcal{G}_{1/\eta_{t}}({\mathbf{u}}^{(k)})\|:=\big\|\eta_{t}^{-1}\cdot\big({\mathbf{u}}^{(k)}-T_{1/\eta_{t}}({\mathbf{u}}^{(k)})\big)\big\|\leq\epsilon_{t}^{\prime},$ (119)
where $T_{1/\eta_{t}}({\mathbf{u}}^{(k)}):=\text{proj}_{\mathcal{X}}\left({\mathbf{u}}^{(k)}-\eta_{t}\nabla F({\mathbf{u}}^{(k)})\right)$, $\eta_{t}$ is a step-size and $\epsilon_{t}^{\prime}$ is a targeted tolerance. Then we take ${\mathbf{x}}^{(t+1)}=T_{1/\eta_{t}}({\mathbf{u}}^{(k)})$ as the approximate solution to (116), (117), or (118). Here, $\mathcal{G}_{1/\eta_{t}}({\mathbf{u}}^{(k)})$ is known as the gradient mapping of $F(\cdot)$ at ${\mathbf{u}}^{(k)}$. By Nesterov [36, Eq. (2.15)], (119) ensures $\text{dist}\left(-\nabla F({\mathbf{x}}^{(t+1)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(t+1)})\right)\leq\epsilon_{t}$ with some $\epsilon_{t}=O(\epsilon_{t}^{\prime})$, i.e., condition (23) holds if $F$ is from (116).
Algorithm 2 Accelerated Projected Gradient (APG) Method [37, Eq. (2.2.63)] for $\min_{{\mathbf{u}}\in\mathcal{X}}F({\mathbf{u}})$
1:Input: targeted tolerance $\epsilon_{t}^{\prime}>0$, step-size $\eta_{t}>0$.
2:Initialization: ${\mathbf{u}}^{(0)}={\mathbf{v}}^{(0)}={\mathbf{x}}^{(t)}$.
3:for iteration $k=0,1,2,\dots$ do
4:  $\tilde{{\mathbf{u}}}^{(k)}=T_{1/\eta_{t}}({\mathbf{u}}^{(k)})$.
5:  if $\|\mathcal{G}_{1/\eta_{t}}({\mathbf{u}}^{(k)})\|=\eta_{t}^{-1}\|{\mathbf{u}}^{(k)}-\tilde{{\mathbf{u}}}^{(k)}\|\leq\epsilon_{t}^{\prime}$ then
6:   Output: ${\mathbf{x}}^{(t+1)}=\tilde{{\mathbf{u}}}^{(k)}$.
7:   Stop.
8:  end if
9:  ${\mathbf{u}}^{(k+1)}=\text{proj}_{\mathcal{X}}({\mathbf{v}}^{(k)}-\eta_{t}\nabla F({\mathbf{v}}^{(k)}))$.
10:  ${\mathbf{v}}^{(k+1)}={\mathbf{u}}^{(k+1)}+\frac{1-\sqrt{q_{F}}}{1+\sqrt{q_{F}}}({\mathbf{u}}^{(k+1)}-{\mathbf{u}}^{(k)})$.
11:end for
In addition, at its $t$-th iteration, SP-LM updates $\mathbf{x}$} by
${\mathbf{x}}^{(t+1)}=\text{proj}_{\mathcal{X}}\left({\mathbf{x}}^{(t)}-\eta_{t}\nabla_{{\mathbf{x}}}\left(f({\mathbf{x}}^{(t)})+{\textstyle\sum_{i=1}^{m}\lambda_{i}^{(t+1)}\cdot g_{i}({\mathbf{x}}^{(t)})}+\frac{p}{2}\|{\mathbf{x}}^{(t)}-{\mathbf{z}}^{(t)}\|^{2}\right)\right),$ (120)
and, at its $t$-th iteration, the SSG method updates $\mathbf{x}$} by
${\mathbf{x}}^{(t+1)}=\begin{cases}\text{proj}_{\mathcal{X}}({\mathbf{x}}^{(t)}-\eta_{t}\nabla f({\mathbf{x}}^{(t)}))\text{ if }\max_{i\in[m]}g_{i}({\mathbf{x}}^{(t)})\leq\epsilon_{t},\\[1.93748pt] \text{proj}_{\mathcal{X}}({\mathbf{x}}^{(t)}-\eta_{t}\boldsymbol{\zeta}_{g}^{(t)})\text{ for some }\boldsymbol{\zeta}_{g}^{(t)}\in\mathrm{conv}(\{\nabla g_{j}({\mathbf{x}}^{(t)}):j\in I({\mathbf{x}}^{(t)})\})\text{ otherwise,}\end{cases}$ (121)
where $I({\mathbf{x}}):=\{j\in[m]:g_{j}({\mathbf{x}})=\max_{i\in[m]}g_{i}({\mathbf{x}})\}$.
In the experiments, we first solve (39) by the projected gradient method up to near optimality to obtain a close approximation of $\mathcal{L}^{*}$ and a solution ${\mathbf{x}}_{\text{feas}}$. Specifically, we use a constant step-size of $0.1$ in the projected gradient method and terminate it when it finds a solution ${\mathbf{x}}_{\text{feas}}$ satisfying $\text{dist}(-\nabla\mathcal{L}({\mathbf{x}}_{\text{feas}}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}_{\text{feas}}))\leq 0.001$. Then we set $\kappa=0.001\mathcal{L}^{*}$ in (41).
The methods in comparison are initialized at ${\mathbf{x}}^{(0)}={\mathbf{x}}_{\text{feas}}$ obtained above. Based on the definitions of $\mathcal{R}$ and $\mathcal{L}$, we can calculate an upper bound of the smooth parameter of the objective and constraint functions in (41) using the data. For the objective function in (41), we notice that $\mathcal{R}({\mathbf{x}})\in[0,1]$ for any ${\mathbf{x}}\in\mathbb{R}^{d}$ and, by a derivation similar to Huang and Lin [16, Eq. (75–77)], $\mathcal{R}({\mathbf{x}})$ and $\nabla\mathcal{R}({\mathbf{x}})$ are Lipschitz continuous with constants of
$\alpha=\frac{1}{4n_{p}}\sum_{i=1}^{n_{p}}\|{\mathbf{a}}_{i}^{p}\|+\frac{1}{4n_{u}}\sum_{i=1}^{n_{u}}\|{\mathbf{a}}_{i}^{u}\|\text{ and }\beta=\frac{1}{4n_{p}}\sum_{i=1}^{n_{p}}\|{\mathbf{a}}_{i}^{p}\|^{2}+\frac{1}{4n_{u}}\sum_{i=1}^{n_{u}}\|{\mathbf{a}}_{i}^{u}\|^{2},$
respectively. Additionally, it is easy to show that $\frac{1}{2}(\mathcal{R}({\mathbf{x}}))^{2}$ is $\alpha$-Lipschitz continuous and $(\beta+\alpha^{2})$-smooth in $\mathbf{x}$}. For the constraint function in (41), by the fact that $\ell(z)$ is 1-Lipschitz continuous and $\sigma(z)$ is $\frac{1}{4}$-Lipschitz continuous, we can show that $\mathcal{L}({\mathbf{x}})$ and $\nabla\mathcal{L}({\mathbf{x}})$ are Lipschitz continuous with constants of
$\gamma=\frac{1}{n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|\text{ and }\gamma^{\prime}=\frac{1}{4n}\sum_{i=1}^{n}\|{\mathbf{a}}_{i}\|^{2},$
respectively. Therefore, (41) satisfies Assumption 3.1 with any $L\geq\max\{\beta+\alpha^{2},\gamma^{\prime}\}$. After estimating $\max\{\beta+\alpha^{2},\gamma^{\prime}\}$ using data, we set $L=10$, 10, and 2.5 on a9a, bank, and COMPAS, respectively, and set $p=2L$. In the APG method applied to (116), (117), and (118), we use a constant step-size $\eta$. On a9a and bank, we select $\eta$ from $\{0.005,0.01,0.02,0.05\}$; on COMPAS, we select $\eta$ from $\{0.02,0.05,0.1,0.2\}$. The choices of other parameters are uniform across all datasets. For the iMELa method, we select $\tau_{t}=\tau$ from $\{5,10,20,50\}$ and $\theta_{t}=\theta$ from $\{0.5,0.75,1\}$, and set $\epsilon_{t}^{\prime}=\frac{c}{t+1}$ with $c$ selected from $\{1,2,5,10\}$. For the iPPP method, following Lin et al. [28, Eq. (5.14)], we set $p_{t}\equiv p$, $\rho_{t}=\rho\sqrt{t+1}$ and $\epsilon_{t}^{\prime}=\frac{1}{\rho_{t}(t+1)}$ with $\rho$ selected from $\{200,500,1000,1500\}$. For DPALM, following Dahal et al. [8, Eq. (21)], we set $\beta_{t}=\beta_{0}\sqrt{t+1}$ and $v_{k}=\frac{v_{0}}{\sqrt{t+1}(\ln(t+1))^{2}}$ with $\beta_{0}$ selected from $\{0.02,0.05,0.1,0.2\}$ on a9a and bank and $\{10^{-4},2\times 10^{-4},5\times 10^{-4},10^{-3}\}$ on COMPAS and $v_{0}$ selected from $\{50,100,150,200\}$; following Dahal et al. [8, Eq. (28) & Sec. 4.5], we set $\epsilon_{t}^{\prime}=\min\{\frac{\epsilon^{\prime}}{8},\frac{1}{2}\sqrt{p/\beta_{t}},1\}$ with $\epsilon^{\prime}:=10^{-2}$ according to Dahal et al. [8, Sec. 4.5]. For SP-LM, we select $\eta$ from $\{0.005,0.01,0.02,0.05\}$ on a9a and bank and from $\{0.02,0.05,0.1,0.2\}$ on COMPAS. We then select $\tau_{t}=\tau$ from $\{5,10,20,50\}$ and $\theta_{t}=\theta$ from $\{0.5,0.75,1\}$. For the SSG method, following Huang and Lin [16], we adopt both static and diminishing step-sizes. For the static step-size, we select $\epsilon_{t}^{\prime}=\epsilon$ from $\{10^{-6},2\times 10^{-6},5\times 10^{-6},10^{-5}\}$ and $\eta_{t}=\eta$ from $\{2\times 10^{-4},5\times 10^{-4},10^{-3},2\times 10^{-3}\}$. For the diminishing step-size, we select $\epsilon_{t}^{\prime}=\frac{E_{1}}{\sqrt{t+1}}$ and $\eta_{t}=\frac{E_{2}}{\sqrt{t+1}}$ and select $E_{1}$ from $\{5\times 10^{-5},10^{-4},2\times 10^{-4},5\times 10^{-4}\}$ and $E_{2}$ from $\{0.02,0.05,0.1,0.2\}$. Since the iMELa method, DPALM, and SP-LM are primal-dual based, and the primal-based iPPP method can employ $\boldsymbol{\lambda}^{(t)}=\rho_{t-1}[{\mathbf{g}}({\mathbf{x}}^{(t)})]_{+}$ as its auxiliary dual estimates according to Step 5 of Algorithm 1 in Lin et al. [28], their best combination of parameters is chosen as the one that minimizes the smallest value of
$\penalty 10000\ \text{dist}\big(-\mathcal{R}({\mathbf{x}}^{(s)})\cdot\nabla\mathcal{R}({\mathbf{x}}^{(s)})-\lambda^{(s)}\cdot\nabla\mathcal{L}({\mathbf{x}}^{(s)}),\mathcal{N}_{\mathcal{X}}({\mathbf{x}}^{(s)})\big)$
$+\big\|\big[\mathcal{L}({\mathbf{x}}^{(s)})-(\mathcal{L}^{*}+\kappa)\big]_{+}\big\|+\big|\lambda^{(s)}\big(\mathcal{L}({\mathbf{x}}^{(s)})-(\mathcal{L}^{*}+\kappa)\big)\big|,$
among iterations $s=0,1,\dots,T_{\text{tuning}}-1$. Here, $T_{\text{tuning}}$ is the number of iterations for parameter tuning. Since the SSG method does not generate $\boldsymbol{\lambda}^{(t)}$, its best combination of parameters is identified as the one that minimizes the smallest value of $\frac{1}{2}(\mathcal{R}({\mathbf{x}}^{(s)}))^{2}$ when $\mathcal{L}({\mathbf{x}}^{(s)})\leq\mathcal{L}^{*}+\kappa+10^{-5}$ among iterations $s=0,1,\dots,T_{\text{tuning}}-1$. For SP-LM and the SSG method, we set $T_{\text{tuning}}=1000$, 5000 and 75000 for a9a, bank and COMPAS, respectively. For a fair comparison, for the iMELa method, the iPPP method, and DPALM, we set $T_{\text{tuning}}$ to be the smallest integer satisfying $\sum_{t=0}^{T_{\text{tuning}}-1}k_{t}\geq 1000\text{, }5000\text{ and }75000$ for a9a, bank and COMPAS, respectively, where $k_{t}$ is the number of inner steps performed at the $t$-th outer iteration. Once all parameters are chosen, we run SP-LM and the SSG method for $T=4T_{\text{tuning}}$ iterations, and run the iMELa method, the iPPP method, and DPALM for $T$ iterations with $T$ being the smallest integer satisfying $\sum_{t=0}^{T-1}k_{t}\geq 4\sum_{t=0}^{T_{\text{tuning}}-1}k_{t}$.